### v1

# Learning to Evolve with Convergence Guarantee via Neural Unrolling

Abstract
The transition from hand-crafted heuristics to data-driven evolutionary algorithms faces a fundamental dilemma: achieving neural plasticity without sacrificing mathematical stability. Emerging learned optimizers demonstrate high adaptability. However, they often lack rigorous convergence guarantees. This deficiency results in unpredictable behaviors on unseen landscapes. To address this challenge, we introduce Learning to Evolve (L2E), a unified bilevel meta-optimization framework. This method reformulates evolutionary search as a Neural Unrolling process grounded in Krasnosel’skii-Mann (KM) fixed-point theory. First, L2E models a coupled dynamic system in which the inner loop enforces a strict contractive trajectory via a structured Mamba-based neural operator. Second, the outer loop optimizes meta-parameters to align the fixed point of the operator with the target objective minimizers. Third, we design a gradient-derived composite solver that adaptively fuses learned evolutionary proposals with proxy gradient steps, thereby harmonizing global exploration with local refinement. Crucially, this formulation provides the learned optimizer with provable convergence guarantees. Extensive experiments demonstrate the scalability of L2E in high-dimensional spaces and its robust zero-shot generalization across synthetic and real-world control tasks. These results confirm that the framework learns a generic optimization manifold that extends beyond specific training distributions.
I Introduction
Evolutionary Algorithms (EAs) constitute a core paradigm in computational intelligence for solving complex black-box optimization problems [1, 2, 3, 4, 5]. These algorithms evolve a population of candidate solutions through selection, recombination, and perturbation. This mechanism enables them to explore highly nonconvex, multimodal, and noisy landscapes without relying on explicit gradient information. Consequently, this paradigm plays a foundational role in computational science and engineering. It underpins diverse applications such as hyperparameter tuning in machine learning [6], aerodynamic shape design [7], molecular conformation prediction [8], and the control of physical systems [9]. Specifically, solvers in these scenarios must navigate high-dimensional and noisy search spaces under tight evaluation budgets. They must simultaneously achieve effective exploration and reliable convergence to high-quality solutions [10].
Optimization research has established two principal families over the past decades: gradient-free, population-based Evolutionary Algorithms (EAs) and model-based Bayesian Optimization (BO). The former family, which includes Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [11], Differential Evolution (DE) [12], and various swarm-based methods [13], is noted for conceptual simplicity and global exploration. Conversely, the latter family is widely adopted for sample efficiency in computationally expensive contexts [14]. However, both families rely on hand-crafted operators and heuristics. Furthermore, they lack a unified theoretical foundation for convergence and stability. This inherent limitation constrains their efficiency in high-dimensional spaces. Consequently, this results in a fundamental lack of adaptability to increasingly complex modern workloads.
A natural response to these limitations is the shift from manually designing operators to data-driven and learnable evolutionary optimizers [15, 16, 17]. Within the Learning to Optimize (L2O) paradigm, recent studies leverage Recurrent Neural Networks (RNNs) and Transformers as controllers, or employ set-based architectures to model optimization trajectories [18]. In parallel, meta-evolution approaches treat mutation-crossover-selection policies as learnable objects. These methods search over expressive families of attention-based, diffusion-inspired, or transformer-style operators [19, 20, 21]. Specifically, they seek to automatically discover evolutionary strategies by learning from a task distribution, such as LQD [22] and the Evolution Transformer [23].
However, a critical gap remains in this transition: the emerging learned optimizers typically face a fundamental dilemma between plasticity and stability. Most existing approaches treat the optimizer as a black-box policy network trained via reinforcement learning to simply minimize trajectory loss. While they demonstrate exceptional plasticity on training tasks, they frequently degenerate into mathematically unconstrained processes. Lacking intrinsic geometric constraints, they are prone to overfitting the training landscape topology, figuratively “memorizing the map” rather than “learning the compass”. Consequently, when deployed on unseen or out-of-distribution tasks, these solvers often exhibit wandering trajectories, catastrophic divergence, or stagnation, failing to provide the stability and rigorous convergence assurances characterizing classical EAs.
Consequently, this limitation motivates us to rethink the fundamental design of learned optimizers: how can we engineer a learned evolutionary optimizer that possesses the adaptive intelligence of neural networks without sacrificing the rigorous convergence laws of classical control? To address this challenge, we propose to ground the design of learned EAs in fixed-point theory. We argue that the fragility of current learned optimizers stems from unconstrained operator learning. A robust evolutionary operator should not merely predict the next step based on heuristics; rather, it must mathematically constitute a contraction mapping. The repeated application of this mapping drives the population towards equilibrium.
Motivated by this insight, we introduce Learning to Evolve (L2E), a unified framework that reformulates evolutionary search as a neural unrolling process. Unlike approaches that view the optimizer merely as a meta-learner, L2E models the optimizer as a learnable dynamical system governed by the classical Krasnosel’skiĭ-Mann (KM) iteration [24]. The core philosophy is to enforce a physics of descent within the neural architecture. In the inner loop, the population update serves as an $\alpha$-averaged operator composed of a learned component and a numerical proxy. This structure enforces non-expansiveness by design. Consequently, the evolution trajectory strictly contracts towards a fixed point, ensuring stability regardless of the landscape’s complexity. In the outer loop, the meta-learner aligns the operator’s fixed point with the target objective minimizers. This alignment effectively shapes the attraction basin of the dynamics. To implement this efficiently in high-dimensional spaces, we propose a structured Mamba-based operator, where spectral normalization constraints are imposed. By leveraging the linear complexity of State Space Models (SSMs), this design overcomes the quadratic bottleneck of Transformers in population-wise interaction complexity, capturing long-range dependencies dynamics efficiently.
This architectural formulation yields the behavioral contrast illustrated in Fig. 1. While conventional learned optimizers often succumb to overfitting the training landscape, which results in unpredictable or unstable trajectories on unseen tasks, L2E induces a stable contraction field where the learned fixed point is explicitly aligned with the target optimum. Consequently, the proposed L2E avoids the pitfall of memorizing the map by learning a generalized descent rule, thereby offering a robust solution for general-purpose adaptation. Our main contributions are:
  • Theory-Driven Formulation: We introduce the Generalized Neural Unrolling (GNU) paradigm. This framework redefines evolution as a curriculum-driven unrolling process under bilevel fixed-point constraints. Furthermore, we establish a joint convergence guarantee to bridge neural plasticity with theoretical optimization rigor.
  • Efficient Composite Solver: We design a Gradient-derived Evolutionary Composite (GEC) solver, which utilizes a Mamba architecture and an elite-mask fusion mechanism. This design addresses the scalability limitations of attention-based optimizers. Consequently, it enables stable meta-evolution in high-dimensional regimes where traditional L2O methods struggle.
  • Rigorous Theoretical Analysis: We provide a fixed-point theoretic analysis establishing the finite-time convergence rate, asymptotic consistency, and bilevel optimality. These guarantees ensure the learned optimizer follows a predictable trajectory, avoiding the instability inherent in unconstrained meta-learning.
The remainder of the paper is organized as follows. Section II reviews related work. Sections III and IV detail the proposed L2E framework and its theoretical analysis. Section V presents the experimental results. Finally, Section VII concludes the paper.
Figure 1: Conceptual contrast. Left: Naive optimizers memorize the map, leading to unstable wandering on unseen landscapes. Right: L2E enforces a physics of descent via non-expansive unrolling, inducing a stable contraction field where the learned fixed-point is aligned with the target optimum.
II Related Work
II-A Learning to Optimize
Learning to Optimize (L2O) blends the prescriptive rigor of classical optimization with the inductive power of machine learning [18]. Specifically, rather than relying on hand-crafted heuristics, L2O seeks to parameterize update rules directly from data. Early works employed model-free architectures, such as RNNs [25] or MLPs [26], to map gradients to updates in a coordinate-wise manner. However, these methods often struggle with scalability and interpretability, primarily due to their opaque nature.
In contrast, model-based L2O approaches embed structural priors into the learner. For instance, Transformer-based optimizers [27] leverage self-attention to capture parameter dependencies. This methodology has further evolved into meta-learning frameworks that train optimizers to adapt across diverse tasks [20, 28]. Notable implementations include reinforcement learning controllers for scheduling [19] and gradient-guided mutation operators [23].
Despite these advances, existing L2O approaches primarily function as unconstrained policy networks. Specifically, they lack intrinsic geometric constraints. This absence implies that there is no theoretical assurance of convergence or stability. Consequently, this deficiency frequently leads to unpredictable behavior when the model transfers to out-of-distribution landscapes, thereby highlighting the necessity for mathematically grounded architectures that enforce stability by design.
II-B Learnable Evolutionary Algorithms
Classical evolutionary algorithms and Bayesian Optimization methods have established the foundation for gradient-free optimization. Prominent examples include Evolutionary Strategies, such as CMA-ES [11], Differential Evolution [12], and L-SHADE [13]. Within the Bayesian domain, notable approaches include SAASBO [29], TuRBO [30], and HEBO [14]. These methods proved successful due to engineered covariance adaptation and population mechanisms [31]. However, their reliance on hand-crafted, non-differentiable rules limits their flexibility in high-dimensional or ill-conditioned landscapes.
To enhance adaptability, recent research has progressively integrated learnability into evolutionary frameworks [32, 33, 34]. Initial efforts utilized surrogates, such as Gaussian Processes, to approximate objective landscapes [35]. Subsequent work has focused on embedding data-driven modules directly into the optimization loop. Examples include meta-evolved operators [36] and context-aware optimizers [19, 20]. Concurrently, researchers have employed Reinforcement Learning for dynamic operator selection [37, 38, 39]. Furthermore, Large Language Models are increasingly being explored for synthesizing heuristic code [40, 41, 42].
Despite this progress, learnable EAs face substantial challenges. Most existing works optimize isolated components rather than the unified evolutionary loop. This isolated optimization results in a fragmented learning process. Furthermore, current methods often lack a consistent mathematical framework to integrate gradient propagation with population search. Consequently, achieving robust zero-shot generalization and scalability in high-dimensional regimes remains a persistent difficulty.
II-C Neural Unrolling
Within the broader inverse problem community, Neural Unrolling (NU) has established itself as a principled methodology for bridging model-based iteration and deep learning. Fundamentally, NU reinterprets an iterative optimization method as a computational graph, wherein each layer corresponds to a single update step and algorithmic parameters function as trainable weights [43]. Although originally developed for signal and image processing, this paradigm has produced influential architectures, including LISTA for sparse coding [44], ADMM-Net for compressive sensing [45], and primal-dual networks for medical image reconstruction [46]. NU embeds domain-specific iterations into differentiable architectures. Recent advances have extended NU to complex visual perception tasks, such as image dehazing and rain removal, by incorporating optimization-inspired heuristics into network design to enhance feature interpretability [47, 48, 49]. Furthermore, in the realm of AutoML, differentiable approaches for Neural Architecture Search (NAS) implicitly leverage the unrolling concept to optimize structural parameters via bilevel gradients [50, 51, 52]. Consequently, it synergizes the interpretability of model-based methods with the adaptability of deep networks, thereby achieving robust performance across diverse inverse problems.
In contrast to its prevalence in signal processing, NU remains underutilized within evolutionary computation. While differentiable EAs exist, they typically rely on ad hoc relaxations rather than systematic unrollings of fixed-point iterations. Thus, the field lacks a framework capable of unifying bilevel meta-learning with rigorous population dynamics guarantees.
II-D Discussion
Existing methodologies navigate a critical trade-off between stability and adaptability. While classical EAs provide theoretical stability, they often lack the adaptability requisite for navigating complex landscapes efficiently. Conversely, emerging L2O approaches offer data-driven flexibility; however, they frequently exhibit instability and poor generalization to out-of-distribution landscapes, primarily due to the absence of intrinsic geometric constraints. To reconcile these diverging paradigms, we propose the L2E that reformulates the evolutionary loop as a neural unrolling process. This formulation is expected to synergize the expressivity of deep networks with the rigorous convergence properties of classical optimization.
Figure 2: Overall framework of the proposed L2E paradigm. Left: Visualizes the interaction where the outer-loop meta-learner reshapes the landscape to align the inner-loop fixed point with the target optimum. Middle: Illustrates the solving process, where the evolutionary search is unrolled into a differentiable pipeline to learn stable trajectories. Right: Details the non-expansive unrolling operator integrates evolutionary and numerical solvers to form a differentiable evolution dynamics.
III Methodology
In this section, we present L2E, a framework designed to unify heuristic search and mathematical optimization. The overall architecture functions as a bilevel dynamical system, as visualized in Fig. 2. To instantiate this paradigm, we first introduce the Generalized Neural Unrolling (GNU) formulation, which provides the theoretical backbone. Second, we detail the Gradient-derived Evolutionary Composite (GEC) solver, which harmonizes gradient-based exploitation with population-based exploration. Finally, we propose a Structured Mamba Neural Operator to efficiently parameterize this process in high-dimensional spaces, thereby ensuring both computational scalability and theoretical stability.
III-A Generalized Neural Unrolling Paradigm
Conventional black-box optimizers often rely on heuristic mechanisms that lack theoretical guarantees. To address this limitation, we reinterpret the optimization process through the lens of non-expansive operator theory. Specifically, we utilize the KM fixed-point formulation [24]. Unlike methods that directly search for optima, the proposed approach optimizes the parameters of a learnable fixed-point operator. This operator implicitly governs the solution trajectory. Formally, let $f:\mathcal{X}\to\mathbb{R}$ denote the black-box objective function. We introduce a learnable non-expansive operator $\mathcal{NU}(\cdot;\bm{\omega})$, parameterized by $\bm{\omega}\in\bm{\Omega}$, whose fixed point approximates the minimizer of $f$. Consequently, we formulate this framework as a constrained bilevel optimization problem [53]:
$\min_{\mathbf{x}\in\mathcal{X},\bm{\omega}\in\Omega}\ell_{\mathtt{meta}}(\mathbf{x};\bm{\omega},\mathtt{Opt}),s.t.\;\mathbf{x}\in\operatorname{Fix}(\mathcal{NU}(\cdot;\bm{\omega})),$ (1)
where $\ell_{\mathtt{meta}}(\mathbf{x};\bm{\omega},\mathtt{Opt})$ denotes a meta-level surrogate objective that encodes task rewards, diversity, or other downstream fitness metrics. Furthermore, $\text{Fix}(\mathcal{NU}):=\{\bm{\mathbf{x}}\mid\bm{\mathbf{x}}=\mathcal{NU}(\bm{\mathbf{x}};\bm{\omega})\}$ defines the set of fixed points. In this formulation, $\bm{\omega}$ represents the outer-level parameters defining the optimization algorithm, whereas $\bm{\mathbf{x}}$ denotes the inner-level population of candidate solutions.
To ensure contractive behavior, we construct $\mathcal{NU}$ as an $\alpha$-averaged operator. Specifically, this formulation employs a convex combination of the identity and a learnable mapping $\mathcal{O}(\cdot,\bm{\omega})$, thereby yielding a KM-style operator:
$\bm{\mathbf{x}}^{k+1}=\mathcal{NU}(\mathbf{x}^{k};\bm{\omega})=(1-\alpha)\cdot\mathbf{x}^{k}+\alpha\cdot\mathcal{O}(\mathbf{x}^{k};\bm{\omega}),$ (2)
$\mathtt{where}\ \mathcal{O}(\cdot;\bm{\omega})\triangleq\{\underbrace{\mathcal{O}_{\mathrm{evo}}(\cdot;\bm{\omega})}_{\mathtt{Neural~operator}}\;\circ\;\underbrace{\mathcal{O}_{\mathrm{num}}(\cdot,\bm{\omega})}_{\mathtt{Numerical~operator}}\}.$ (3)
The operator $\mathcal{NU}(\cdot;\bm{\omega})$ is defined as a composition of two components, modulated by an update relaxation parameter $\alpha\in(0,1]$. Specifically, $\mathcal{O}_{\text{num}}$ denotes a numerical solver, such as smoothing or projection, while $\mathcal{O}_{\text{evo}}$ encapsulates a learnable search policy. By enforcing non-expansiveness on $\mathcal{O}$, the GNU paradigm ensures that the evolution trajectory $\mathbf{x}^{0}\to\dots\to\mathbf{x}^{K}$ follows a strictly contractive path. Consequently, this formulation provides a theoretical safeguard against divergence, which addresses a common failure mode in standard learned optimizers.
This formulation gives rise to a semantically structured learning process: i) Inner-Level (Evolve): For fixed algorithm parameters $\bm{\mathbf{\omega}}$, the inner loop performs K-step unrolled evolution of an initial population $\bm{\mathbf{x}}^{0}$. This unrolling implicitly defines the optimization trajectory, transforming initial samples into refined candidates. ii) Outer-Level (Meta-Train): The outer loop treats the terminal population $\bm{\mathbf{x}}^{K}$ as a response to the optimizer parameters $\bm{\omega}$. A meta-objective $\ell_{\mathtt{meta}}(\bm{\mathbf{x}}^{K};\bm{\omega})$ is computed to measure task performance, and gradients are backpropagated through the unrolling path to update $\bm{\omega}$: $\bm{\omega}^{t+1}=\bm{\omega}^{t}-\gamma\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}(\bm{\mathbf{x}}^{K};\bm{\omega}^{t}),$ where $\gamma$ is the meta-step size. This teaches the operator to evolve more effective trajectories across tasks.
1
Input : Initial population $\mathbf{x}_{0}\in\mathcal{X}$, Meta-parameters $\bm{\omega}$, Learning rate $\gamma$, Unroll horizon $K$, Outer iterations $T$, Relaxation factor $\alpha\in(0,1]$, Differentiable operators $\mathcal{O}_{\mathrm{evo}}$ (Mamba) / $\mathcal{O}_{\mathrm{num}}$ (Proxy)
Output : Optimized population $\mathbf{x}^{K}$, meta-parameters $\bm{\omega}^{T}$
2
3 Initialization: $\bm{\omega}^{0}\leftarrow\bm{\omega}_{\text{init}}$;
4for $t=0$ to T1T-1 do
  $\mathbf{x}^{0}\leftarrow\mathbf{x}_{0}$
  // Reset Trajectory
5
  /* Differentiable Evolution Unrolling (Inner-Level) */
6  for $k=1$ to $K$ do
     /* Step 1: Non-expansive Neural Evolution (KM Iteration) */
7     $\mathbf{d}_{IL}^{k}\leftarrow\mathbf{x}^{k-1}+\alpha\cdot\left(\mathcal{O}_{\mathrm{evo}}(\mathbf{x}^{k-1},\bm{\omega})-\mathbf{x}^{k-1}\right)$
8
    /* Step 2: Numerical Guidance */
9     if is training then
        $\mathbf{g}_{\text{proxy}}\leftarrow\bm{P}_{\bm{\omega}}^{-1}\nabla_{\mathbf{x}}\ell_{\mathtt{meta}}(\mathbf{x}^{k-1};\bm{\omega}^{t})$
        // Gradient Injection
10        $\mathbf{d}_{OL}^{k}\leftarrow\mathbf{x}^{k-1}-s_{k}\cdot\mathbf{g}_{\text{proxy}}$
11
12    else
        $\mathbf{d}_{OL}^{k}\leftarrow\mathbf{x}^{k-1}$
        // Evaluation Mode
13
14
    /* Step 3: Adaptive Manifold Rectification */
     $\Delta f^{k}\leftarrow f(\mathbf{d}_{OL}^{k})-f(\mathbf{d}_{IL}^{k})$;    $\bm{M}^{k}\leftarrow\sigma(-\Delta f^{k}/\tau)$
     // Soft Gating
15     $\mathbf{x}^{k}\leftarrow\text{Proj}_{\mathcal{X},\bm{P}_{\bm{\omega}}}\left(\bm{M}^{k}\odot\bm{d}_{OL}^{k}+(1-\bm{M}^{k})\odot\bm{d}_{IL}^{k}\right)$
16
  /* Meta-Update via BPTT (Outer-Level) */
  $\ell_{\mathtt{meta}}(\bm{\omega}^{t})\leftarrow-\mathbb{E}_{f\sim p(f)}\left[\frac{\bar{f}(\mathbf{x}^{0})-\bar{f}(\mathbf{x}^{K}(\bm{\omega}^{t}))}{|\bar{f}(\mathbf{x}^{0})|+\epsilon}\right]$
  // Cumulative Regret
17
18 $\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}\leftarrow\textnormal{{{AutoDiff}}}_{\text{BPTT}}(\ell_{\mathtt{meta}},\{\mathbf{x}^{k}\}_{k=0}^{K})$
  $\bm{\omega}^{t+1}\leftarrow\bm{\omega}^{t}-\gamma\cdot\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}$
  // Meta-Update
19
20
21Return $\mathbf{x}^{K}$, $\bm{\omega}^{T}$
Algorithm 1 L2E: Differentiable Bilevel Evolution Strategy. The algorithm couples a neural fixed-point iteration with a dual-mode numerical operator: proxy gradients are injected as search directions during training, while inference remains purely evaluation-based.
III-B Gradient-Derived Evolutionary Composite Solver
To instantiate the non-expansive operator $\mathcal{NU}(\cdot;\bm{\omega})$ introduced above, we propose a hybrid mechanism that fuses the global exploration capability of evolutionary algorithms with the local refinement power of gradient descent. This Gradient-Derived Evolutionary Composite (GEC) solver operates in a dual-mode fashion.
III-B1 Learned evolutionary update
At each inner iteration $k$, given the current population stat $\bm{\mathbf{x}}^{k-1}$, the solver computes two candidate update directions:
$\bm{d}_{IL}^{k}(\bm{\omega})=\mathcal{NU}(\bm{\mathbf{x}}^{k-1};\bm{\omega}),$ (4)
where $\mathcal{NU}$ follows the KM-style fixed-point update defined in Eq. (2), composed of the learnable operator $\mathcal{O}_{\bm{\omega}}$.
III-B2 Proxy gradient update
$\bm{d}_{OL}^{k}(\bm{\omega})=\bm{\mathbf{x}}^{k-1}-s_{k}\,\bm{P}_{\bm{\omega}}^{-1}\nabla_{\bm{\mathbf{x}}}\,\ell_{\mathtt{meta}}(\bm{\mathbf{x}}^{k-1},\bm{\omega}),$ (5)
where $s_{k}$ denotes the inner step size and $\bm{P}_{\bm{\omega}}$ is a preconditioning matrix (e.g., identity, diagonal Hessian, or low-rank approximation).
III-B3 Composite update with soft-gated fusion
To seamlessly integrate these paths, we introduce an Adaptive Manifold Rectification mechanism. A dimension-wise gating mask $\mathbf{M}^{k}$ is computed based on the relative fitness improvement $\Delta f^{k}=f(\bm{d}_{OL}^{k})-f(\bm{d}_{IL}^{k})$, and apply a smooth gating function to produce a soft selection mask
$\bm{M}^{k}=\sigma(-\Delta f^{k}/\tau),\quad\bm{M}^{k}\in(0,1)^{B\times N\times 1},$ (6)
where $\sigma(\cdot)$ is the element-wise sigmoid function and $\tau$ is a temperature parameter controlling gating sharpness. The final update is a soft interpolation between the two candidates:
$\bm{\mathbf{x}}^{k}=\text{Proj}_{\mathcal{X},\bm{P}_{\bm{\omega}}}\left(\bm{M}^{k}\odot\bm{d}_{OL}^{k}+(1-\bm{M}^{k})\odot\bm{d}_{IL}^{k}\right),$ (7)
where $\text{Proj}_{\mathcal{X},\bm{P}_{\bm{\omega}}}$ denotes a projection back to the feasible domain under the $\bm{P}_{\bm{\omega}}$-induced metric. This fusion mechanism balances exploitation and exploration in a differentiable, stability-preserving manner.
III-B4 Meta-Objective Function
To ensure robust meta-optimization across varying problem scales, we define $\ell_{\mathtt{meta}}$ based on the relative reduction of objective values. Let $\bar{f}(\mathbf{x})$ denote the mean objective value of a population $\mathbf{x}$. The meta-objective is formulated to minimize the negative expected normalized improvement over the task distribution $p(f)$:
$\ell_{\mathtt{meta}}(\bm{\omega})=-\mathbb{E}_{f\sim p(f)}\left[\frac{\bar{f}(\mathbf{x}^{0})-\bar{f}(\mathbf{x}^{K}(\bm{\omega}))}{|\bar{f}(\mathbf{x}^{0})|+\epsilon}\right],$ (8)
where $\mathbf{x}^{K}(\bm{\omega})$ is the terminal population evolved on function $f$, and $\epsilon$ is a small constant for numerical stability. By minimizing this normalized loss, the framework is encouraged to maximize the proportion of cost reduction achieved within the unrolling horizon $K$, ensuring consistent optimization efficiency regardless of the absolute magnitude of the initial objective values. The complete execution protocol is formalized in Algorithm 1. This pseudocode unifies the theoretical framework into a coherent workflow: The inner loop (Lines 4-12) executes the composite update to enforce strict population contraction, while the outer loop (Lines 13-15) performs meta-optimization via BackPropagation Through Time (BPTT). By minimizing the derived meta-loss (Eq. 8), the algorithm effectively aligns the operator’s fixed point with the target optimum, bridging the gap between theoretical design and practical implementation.
III-C Structured Mamba Neural Operator
To instantiate the evolution operator $\mathcal{O}_{evo}(\cdot;\omega)$, we introduce the Mamba-enhanced evolutionary operator, as illstrated in Fig. 3. Unlike sequential SSMs, we implement this as a parallelized selective block to accommodate the permutation invariance of evolutionary populations, fusing local state refinement with global exploration.
III-C1 Parallel Mamba Block
The core of this operator lies in the selection mechanism, where the SSM parameters are derived directly from the input to filter noise and propagate promising signals. We transform the raw population $\mathbf{x}\in\mathbb{R}^{B\times N\times D}$ and corresponding fitness values $\mathbf{f}$ into a compact embedding $\mathbf{E}$. Then we project $\mathbf{E}$ to generate the input-dependent discretization timescale $\bm{\Delta}$ and state transformation matrices $\mathbf{B},\mathbf{C}$. We decompose the operator into two parallel streams, a state expansion stream and an input gating stream, to approximate the discretized SSM dynamics without sequential scanning. The latent state $\mathbf{s}$ and the passthrough signal $\mathbf{u}$ (i.e., $\mathbf{u}=\phi_{in}(\mathbf{E})$) are computed as:
$M_{\mathbf{s}}=(\bm{\Delta}\mathbf{B})\odot\mathbf{E},\,\,[\bm{\Delta},\mathbf{B},\mathbf{C}]=\text{Linear}_{proj}(\mathbf{E}).$ (9)
Here, $M_{\mathbf{s}}$ represents the transformed evolutionary state modulated by the selective timescale $\bm{\Delta}$. The final output of the Mamba block is regulated by a learnable multiplicative gate $\mathbf{z}$, enabling dynamic information retention:
$\small\mathbf{H}_{mamba}=\text{LN}\left(\mathbf{z}\odot M_{\mathbf{s}}+(1-\mathbf{z})\odot\mathbf{u}\right),\mathbf{z}=\sigma(\text{Linear}_{z}(\mathbf{E})).$ (10)
This gated formulation ($\mathbf{z}\odot M_{\mathbf{s}}+(1-\mathbf{z})\odot\mathbf{u}$) mathematically mirrors the selective retention mechanism of Gated SSMs while ensuring stable gradient flow.
III-C2 Hybrid Adaptive Fusion
To complement the local refinement of Mamba with global population-level coordination, we integrate a parallel Multi-Head Self-Attention (MHSA) path: $\mathbf{H}_{attn}=\text{MHSA}(\mathbf{E})+\mathbf{E}.$ The final evolutionary proposal $\Delta_{evo}$ is synthesized via an adaptive routing mechanism that dynamically balances state exploitation and global exploration. The routing weights are determined by a lightweight non-linear router that processes population-level statistics. Specifically, a feature vector $\mathbf{s}_{pop}\in\mathbb{R}^{3}$ containing fitness standard deviation, fitness range, and population diversity is fed into an MLP to generate logits, which are then normalized:
$\lambda_{ssm},\lambda_{attn}]=\text{Softmax}(\text{MLP}_{router}(\mathbf{s}_{pop})).$ (11)
These adaptive coefficients then weigh the projected outputs of the two paths to form the final update proposal:
$\small\Delta_{evo}=\lambda_{ssm}\cdot\text{Tanh}(\Psi_{m}(\mathbf{H}_{mamba}))+\lambda_{attn}\cdot\text{Tanh}(\Psi_{a}(\mathbf{H}_{attn})).$ (12)
where $\Psi_{m}$ and $\Psi_{a}$ are projection heads, and the Tanh activation ensures the proposed updates are bounded within a controlled range for stability. Crucially, to strictly satisfy the theoretical Lipschitz continuity requirement of the GNU framework, we apply Spectral Normalization to all learnable linear transformations within both the Mamba and Attention blocks. By constraining the spectral norm of weight matrices, we ensure that the global Lipschitz constant of the operator $\mathcal{O}_{evo}$ is bounded (effectively $\leq 1$), thereby guaranteeing the contractive stability of the unrolled fixed-point iteration.
Figure 3: Architecture of the Structured Mamba Neural Operator.
IV Theoretical Analysis
In this section, we establish the theoretical foundation of L2E. Specifically, we demonstrate that L2E satisfies convergence theorems under a set of regularity assumptions. In contrast to heuristic meta-optimizers, which typically lack stability guarantees, the proposed non-expansive unrolling mechanism ensures a predictable convergence trajectory. The analysis proceeds in hierarchical steps. First, we quantify the finite-time convergence rates. Second, we establish the asymptotic consistency and bilevel optimality of the framework under stronger topological conditions. Due to space constraints, we provide detailed proofs for all propositions and theorems in the Supplementary Document.
IV-A Convergence Rate Analysis
To characterize the convergence rate, we first adopt standard regularity assumptions gritounded in KM theory.
Assumption IV.1.
  1. (A1)
    (Non-expansiveness) For any fixed $\bm{\omega}\in\Omega$, the learned composite operator $\mathcal{O}(\cdot;\bm{\omega})$ is non-expansive with respect to the population state $\mathbf{x}$: $\|\mathcal{O}(\mathbf{x};\bm{\omega})-\mathcal{O}(\mathbf{x}^{\prime};\bm{\omega})\|\leq\|\mathbf{x}-\mathbf{x}^{\prime}\|,\quad\forall\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}.$
  2. (A2)
    (Smoothness) $\ell_{\mathtt{meta}}(\mathbf{x},\bm{\omega})$ is $L_{\mathbf{x}}$-smooth w.r.t. the population $\mathbf{x}$ and $L_{\bm{\omega}}$-smooth w.r.t. the parameters $\bm{\omega}$.
  3. (A3)
    (Lipschitz Jacobian) The unrolled mapping $\mathbf{x}^{K}(\bm{\omega})$ admits a bounded Jacobian, i.e., $\|\nabla_{\bm{\omega}}\mathbf{x}^{K}(\bm{\omega})\|\leq C_{K}$.
  4. (A4)
    (Bounded domain) The initial population $\mathbf{x}^{0}$ and $\mathbf{x}^{*}(\bm{\omega})\in\text{Fix}(\mathcal{NU}(\cdot;\bm{\omega}))$ reside within a bounded region of diameter $D_{0}$: $\|\mathbf{x}^{0}-\mathbf{x}^{*}(\bm{\omega})\|\leq D_{0}$.
Based on these, we derive the approximation error of the inner loop and the convergence rate of the outer loop.
Proposition IV.1 (Approximation Error).
Under Assumption (A1), the $K$-step unrolled KM iteration satisfies:
$\|\mathbf{x}^{K}(\bm{\omega})-\mathbf{x}^{*}(\bm{\omega})\|\leq(1-\alpha)^{K}D_{0},$ (13)
where $\mathbf{x}^{*}(\bm{\omega})$ is the fixed point satisfying $\mathbf{x}^{*}=\mathcal{O}(\mathbf{x}^{*};\bm{\omega})$.
Theorem IV.1 (Convergence Rate).
Under Assumptions (A1)-(A4), let the meta-parameters be updated via gradient descent: $\bm{\omega}^{t+1}=\bm{\omega}^{t}-\gamma\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}(\mathbf{x}^{K}(\bm{\omega}^{t});\bm{\omega}^{t})$. With step size $\gamma<\frac{1}{L_{\bm{\omega}}+L_{\mathbf{x}}C_{K}}$, the meta-gradient norm satisfies:
$\min_{0\leq t<T}\|\nabla\ell_{\mathtt{meta}}(\bm{\omega}^{t})\|^{2}\leq\mathcal{O}\left(\frac{1}{T}+(1-\alpha)^{2K}\right).$ (14)
Remark: The bound consists of an optimization error $\mathcal{O}(1/T)$ and an approximation bias $\mathcal{O}((1-\alpha)^{2K})$, confirming that L2E consistently improves the operator towards a stationary point of the meta-loss.
IV-B Asymptotic Consistency and Bilevel Optimality
To analyze the limiting behavior of the composite solver (Algorithm 1) and the exactness of the bilevel solution, we introduce stronger topological assumptions regarding the metric space induced by the preconditioner $\mathbf{P}_{\bm{\omega}}$.
Assumption IV.2.
  1. (B1)
    (Non-expansiveness) For each $\bm{\omega}\in\Omega$, the composite operator satisfies: $\|\mathcal{O}(\mathbf{u}_{1};\bm{\omega})-\mathcal{O}(\mathbf{u}_{2};\bm{\omega})\|_{\mathbf{P}_{\bm{\omega}}}\leq\|\mathbf{u}_{1}-\mathbf{u}_{2}\|_{\mathbf{P}_{\bm{\omega}}},\quad\forall\mathbf{u}_{1},\mathbf{u}_{2}\in\mathcal{X}.$
  2. (B2)
    (Closedness) The graph $\mathrm{gph}(\mathcal{O}(\cdot;\bm{\omega})):=\{(\mathbf{u},\mathbf{v})\in\mathcal{X}\times\mathcal{X}\mid\mathbf{v}=\mathcal{O}(\mathbf{u};\bm{\omega})\}$, is closed.
  3. (B3)
    (Compactness and Continuity) The sets $\Omega$ and $\mathcal{X}$ are compact. For any $\bm{\omega}$, $\text{Fix}(\mathcal{NU}(\cdot;\bm{\omega}))$ is non-empty. $\ell_{\mathtt{meta}}$ is continuous on $\mathcal{X}\times\Omega$, and for fixed $\bm{\omega}$, $\ell_{\mathtt{meta}}(\cdot;\bm{\omega})$ is $L_{\ell}$-smooth, convex, and lower-bounded.
Under these structural conditions, we establish two fundamental properties: the asymptotic consistency of the inner solver and the optimality of the bilevel solution.
Theorem IV.2.
Under Assumptions IV.2, we establish the following convergence properties for the L2E framework:
(i) Asymptotic Consistency: Under Assumptions IV.2, let $\{\mathbf{x}^{k}(\bm{\omega})\}$ be the sequence generated by Algorithm 1. For any $\bm{\omega}\in\Omega$, provided the step size satisfies $s_{k}=\frac{\kappa}{k+1}$ with $0<\kappa<\frac{\lambda_{\min}(\mathbf{P}_{\bm{\omega}})}{L_{\mathbf{x}}}$, the sequence converges to the fixed-point set, i.e., $\lim_{k\to\infty}\mathrm{dist}(\mathbf{x}^{k}(\bm{\omega}),\mathrm{Fix}(\mathcal{NU}(\cdot;\bm{\omega})))=0$, and the meta-objective value converges to the optimum:
$\lim_{k\to\infty}\ell_{\mathtt{meta}}(\mathbf{x}^{k}(\bm{\omega});\bm{\omega})=\inf_{\mathbf{x}\in\mathrm{Fix}(\mathcal{NU}(\cdot;\bm{\omega}))}\ell_{\mathtt{meta}}(\mathbf{x};\bm{\omega}).$ (15)
(ii) Bilevel Optimality: As the unrolling depth $K\to\infty$, let $\{(\mathbf{x}^{K}(\bm{\omega}^{K}),\bm{\omega}^{K})\}$ be the sequence of generated solutions. We prove that any limit point $(\bar{\mathbf{x}},\bar{\bm{\omega}})$ constitutes a valid solution to the fixed-point constrained bilevel problem (Eq. 1), satisfying $\bar{\bm{\omega}}\in\arg\min_{\bm{\omega}\in\Omega}\varphi(\bm{\omega})$ and $\bar{\mathbf{x}}=\mathcal{NU}(\bar{\mathbf{x}};\bar{\bm{\omega}})$. Furthermore, the sequence of surrogate objectives converges to the true optimum, i.e., $\lim_{K\to\infty}\inf_{\bm{\omega}\in\Omega}\varphi_{K}(\bm{\omega})=\inf_{\bm{\omega}\in\Omega}\varphi(\bm{\omega})$.
V Experimental Study
V-A Experimental Setting
Benchmark Setup. We evaluate the optimization efficiency and zero-shot generalization capability of L2E using a hierarchical benchmark suite. This suite ranges from synthetic functions to real-world robot control tasks. We structure the evaluation protocol into three categories based on the relationship between the training and testing distributions:
  • Synthetic Benchmarks (In-Distribution): We utilize BBOB-10D, BBOB-30D [54], and LSGO-1000D [55] as the primary testbeds. These datasets constitute both the training sets and the in-distribution test sets. Specifically, the BBOB tasks evaluate basic optimization capabilities with budgets of $2\times 10^{4}$ (10D) and $5\times 10^{4}$ (30D). In contrast, LSGO assesses scalability in high-dimensional regimes ($D\approx 1000$) with a budget of $3\times 10^{6}$.
  • Out-of-Distribution (OOD) Generalization: We select two benchmarks to assess the transferability of L2E across disparate domains. First, we employ the UAV benchmark [56], which comprises 56 real-world path planning instances. Here, we test a model trained exclusively on BBOB-30D under a constrained budget of $2.5\times 10^{2}$. Second, we utilize BBOB-surrogate-10D, i.e., a synthetic variant derived from surrogate model landscapes. This benchmark tests robustness against landscape shifts and approximation errors. In this case, we apply the model trained on standard BBOB-10D directly without fine-tuning.
  • Neuroevolution for Robot Control (Flexibility Verification): We extend the evaluation to the NE benchmark [57], which consists of high-dimensional robot control policy search tasks ($D>1000$). We use this suite specifically to benchmark L2E against reinforcement learning methods. This comparison assesses the flexibility of the framework in control scenarios, given a maximum evaluation budget of $2,500$.
TABLE I: Numerical performance evaluation (mean $\pm$ std) of representative algorithms on the BBOB-10D benchmark. Each entry is calculated from 10 independent test runs. The best and second-best results are highlighted using bold and bold, respectively.
Method Buche_Ras Attractive_Sec Step_Ell Rosenbrock_ori Rosenbrock_rot Ellipsoidal_hc Discus Bent_Cig
PSO 3.429E+02 1.109E+03 8.667E+01 4.003E+03 1.798E+03 3.288E+05 7.338E+01 1.768E+07
($\pm$6.591E+01) ($\pm$7.005E+02) ($\pm$1.083E+01) ($\pm$1.471E+03) ($\pm$7.772E+02) ($\pm$2.522E+05) ($\pm$3.266E+01) ($\pm$1.077E+07)
DE 1.012E+02 4.934E+01 3.099E+01 3.527E+02 7.201E+02 2.859E+04 4.860E+01 2.070E+06
($\pm$2.988E+01) ($\pm$6.491E+00) ($\pm$7.657E+00) ($\pm$1.685E+02) ($\pm$4.851E+02) ($\pm$2.196E+04) ($\pm$6.461E+00) ($\pm$5.398E+05)
CMAES 1.514E+03 1.838E+04 8.980E+01 1.831E+04 2.004E+02 1.612E+06 6.288E+01 7.855E+07
($\pm$4.183E+00) ($\pm$1.386E+02) ($\pm$3.532E-01) ($\pm$3.221E+02) ($\pm$6.800E+00) ($\pm$6.243E+03) ($\pm$1.963E-01) ($\pm$3.370E+05)
SAHLPSO 1.424E+02 1.167E+02 2.402E+01 9.622E+02 1.070E+03 4.344E+04 1.141E+02 7.618E+06
($\pm$6.008E+00) ($\pm$8.459E+01) ($\pm$1.786E+01) ($\pm$6.640E+02) ($\pm$6.981E+02) ($\pm$2.739E+04) ($\pm$2.008E+01) ($\pm$2.039E+06)
RNNOPT 6.616E+03 6.720E+04 7.170E+01 1.930E+04 6.619E+01 3.805E+06 1.740E+05 6.700E+07
($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00)
DEDQN 3.627E+02 9.099E+03 1.018E+02 1.369E+04 1.021E+04 5.756E+05 2.154E+02 3.578E+07
($\pm$9.568E+01) ($\pm$1.020E+04) ($\pm$2.562E+01) ($\pm$3.021E+03) ($\pm$3.073E+03) ($\pm$1.608E+05) ($\pm$5.994E+01) ($\pm$1.490E+07)
LES 1.473E+03 1.816E+03 3.076E+02 2.697E+03 1.334E+03 3.207E+05 9.363E+02 1.308E+07
($\pm$4.452E+01) ($\pm$1.624E+03) ($\pm$7.422E-01) ($\pm$2.109E+02) ($\pm$4.173E+00) ($\pm$6.588E+04) ($\pm$8.504E-01) ($\pm$1.848E+06)
LGA 2.462E+02 4.033E+02 1.794E+02 5.188E+03 3.062E+04 2.578E+05 2.265E+02 5.390E+07
($\pm$5.814E+01) ($\pm$0.000E+00) ($\pm$9.467E+00) ($\pm$2.615E+03) ($\pm$1.402E+04) ($\pm$8.073E+04) ($\pm$7.462E+01) ($\pm$1.930E+07)
GLHF 4.513E+02 1.689E+04 9.959E+01 1.324E+04 1.156E+04 4.462E+05 7.396E+02 3.449E+07
($\pm$1.747E+02) ($\pm$1.254E+04) ($\pm$1.196E+01) ($\pm$3.514E+03) ($\pm$2.001E+03) ($\pm$1.668E+05) ($\pm$5.473E+02) ($\pm$3.621E+06)
B2OPT 2.734E+02 1.116E+02 1.061E+01 3.067E+02 2.422E+01 3.050E+04 4.083E+01 7.585E+06
($\pm$4.075E+01) ($\pm$4.417E+00) ($\pm$3.472E+00) ($\pm$2.971E+02) ($\pm$3.558E+00) ($\pm$1.391E+04) ($\pm$3.942E+00) ($\pm$1.256E+06)
[…]
Method Sharp_Rid Different_Pow Schaffers_hig Composite_Gri Schwefel Gallagher_21P Katsuura Lunacek_bi_Ras
PSO 9.431E+02 7.173E+00 2.411E+01 4.266E+00 3.235E+02 6.513E+01 1.209E+00 1.166E+02
($\pm$1.627E+02) ($\pm$3.197E+00) ($\pm$9.562E-01) ($\pm$6.784E-01) ($\pm$3.334E+02) ($\pm$5.339E+00) ($\pm$2.412E-01) ($\pm$7.371E+00)
DE 2.583E+02 2.088E+00 1.081E+01 4.509E+00 2.413E+00 1.420E+01 2.733E+00 8.125E+01
($\pm$2.861E+01) ($\pm$1.005E+00) ($\pm$8.182E-01) ($\pm$2.996E-01) ($\pm$9.219E-02) ($\pm$1.058E+01) ($\pm$7.799E-01) ($\pm$3.929E+00)
CMAES 1.309E+03 1.309E+01 4.228E+01 3.101E+00 1.671E+03 7.707E+01 5.306E-01 6.387E+01
($\pm$1.568E+00) ($\pm$4.955E-02) ($\pm$5.127E-03) ($\pm$5.885E-01) ($\pm$4.225E+01) ($\pm$1.659E-02) ($\pm$6.150E-02) ($\pm$4.391E-02)
SAHLPSO 3.679E+02 1.316E+00 1.229E+01 5.759E+00 3.942E+01 1.675E+01 1.577E+00 9.341E+01
($\pm$2.165E+01) ($\pm$1.267E+00) ($\pm$5.473E+00) ($\pm$5.528E-01) ($\pm$3.132E+01) ($\pm$5.020E+00) ($\pm$5.575E-01) ($\pm$2.748E+00)
RNNOPT 1.798E+03 1.069E+01 5.223E+01 2.056E+00 8.060E+03 8.486E+01 3.440E+00 1.221E+02
($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00)
DEDQN 1.057E+03 1.047E+01 3.073E+01 1.127E+01 4.618E+03 5.834E+01 3.613E+00 1.480E+02
($\pm$1.827E+02) ($\pm$3.672E+00) ($\pm$4.258E+00) ($\pm$1.689E+00) ($\pm$8.773E+02) ($\pm$9.502E+00) ($\pm$5.164E-01) ($\pm$2.998E+01)
LES 3.274E+03 1.505E+03 7.211E+02 1.700E+03 2.104E+03 2.454E+03 6.017E+02 1.270E+03
($\pm$6.658E+01) ($\pm$3.239E-01) ($\pm$6.131E-01) ($\pm$3.196E-02) ($\pm$2.379E-01) ($\pm$1.005E+01) ($\pm$1.436E-01) ($\pm$2.917E+00)
LGA 9.683E+02 1.663E+01 2.613E+01 1.322E+01 1.234E+04 4.875E+01 1.618E+00 1.778E+02
($\pm$2.034E+02) ($\pm$5.850E+00) ($\pm$4.232E+00) ($\pm$5.755E+00) ($\pm$3.810E+03) ($\pm$1.666E+01) ($\pm$0.000E+00) ($\pm$1.977E+01)
GLHF 1.111E+03 1.530E+01 2.792E+01 9.223E+00 4.686E+03 4.876E+01 3.299E+00 1.499E+02
($\pm$2.052E+02) ($\pm$8.705E-01) ($\pm$3.296E+00) ($\pm$1.221E+00) ($\pm$1.331E+03) ($\pm$1.684E+01) ($\pm$4.095E-01) ($\pm$2.572E+01)
B2OPT 4.721E+02 4.665E+00 1.176E+01 1.633E+00 2.877E+00 2.987E+01 2.293E+00 7.342E+01
($\pm$6.671E+01) ($\pm$1.035E+00) ($\pm$2.675E+00) ($\pm$1.135E-01) ($\pm$3.331E-01) ($\pm$1.920E+01) ($\pm$4.472E-02) ($\pm$5.952E+00)
[…]
indicates ours variant enhanced with the proposed Mamba-based neural operator.
TABLE II: Numerical performance evaluation (mean $\pm$ std) of representative algorithms on the BBOB-30D benchmark.. Each entry is calculated from 10 independent test runs. The best and second-best results are highlighted using bold and bold, respectively.
Method Buche_Ras Attractive_Sec Step_Ell Rosenbrock_ori Rosenbrock_rot Ellipsoidal_hc Discus Bent_Cig
PSO 2.169E+03 1.224E+05 7.002E+02 1.109E+04 5.884E+02 1.110E+06 1.633E+02 1.159E+08
($\pm$5.998E+02) ($\pm$5.208E+04) ($\pm$4.272E+01) ($\pm$2.664E+03) ($\pm$2.849E+02) ($\pm$2.838E+05) ($\pm$2.059E+01) ($\pm$1.495E+07)
DE 4.996E+02 8.765E+02 1.224E+02 2.471E+03 2.814E+03 3.085E+05 1.820E+02 2.201E+07
($\pm$1.715E+01) ($\pm$7.413E+02) ($\pm$3.713E+01) ($\pm$7.355E+02) ($\pm$5.186E+02) ($\pm$1.613E+05) ($\pm$1.608E+01) ($\pm$3.474E+06)
CMAES 4.236E+03 1.409E+05 1.435E+03 4.085E+04 1.850E+02 2.423E+06 1.732E+02 1.689E+08
($\pm$4.880E+01) ($\pm$1.602E+03) ($\pm$1.935E+01) ($\pm$1.404E+02) ($\pm$3.947E+00) ($\pm$1.120E+04) ($\pm$6.706E-01) ($\pm$1.487E+06)
SAHLPSO 9.574E+02 5.158E+04 4.115E+02 2.687E+04 9.397E+03 1.296E+06 2.760E+02 1.124E+08
($\pm$6.404E+01) ($\pm$4.482E+04) ($\pm$1.473E+02) ($\pm$1.173E+04) ($\pm$2.132E+03) ($\pm$5.424E+05) ($\pm$7.851E+01) ($\pm$1.062E+08)
RNNOPT 1.603E+04 5.768E+05 1.511E+03 9.192E+04 1.975E+02 1.431E+07 4.950E+06 4.499E+08
($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$1.455E-11) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00)
DEDQN 2.061E+03 2.325E+05 9.007E+02 1.712E+05 1.518E+05 3.772E+06 5.374E+02 2.796E+08
($\pm$3.744E+02) ($\pm$4.669E+04) ($\pm$4.917E+01) ($\pm$1.802E+04) ($\pm$9.146E+03) ($\pm$4.637E+05) ($\pm$2.222E+02) ($\pm$3.171E+07)
LES 5.822E+03 3.620E+05 3.044E+03 5.335E+04 1.485E+03 3.974E+06 1.953E+03 2.468E+08
($\pm$3.715E+02) ($\pm$2.230E+04) ($\pm$1.385E+02) ($\pm$9.407E+02) ($\pm$5.223E-01) ($\pm$2.115E+05) ($\pm$3.210E+01) ($\pm$1.006E+07)
LGA 1.082E+04 6.709E+05 1.266E+03 2.607E+05 2.086E+05 6.214E+06 1.187E+05 4.517E+08
($\pm$4.568E+03) ($\pm$2.485E+05) ($\pm$1.630E+02) ($\pm$6.975E+04) ($\pm$6.213E+04) ($\pm$2.719E+06) ($\pm$1.669E+05) ($\pm$9.124E+07)
GLHF 4.510E+03 3.470E+05 1.275E+03 3.373E+05 1.993E+05 7.354E+06 1.892E+03 5.111E+08
($\pm$5.042E+02) ($\pm$2.271E+05) ($\pm$3.484E+02) ($\pm$2.066E+04) ($\pm$2.834E+04) ($\pm$1.035E+06) ($\pm$1.603E+03) ($\pm$6.690E+07)
B2OPT 1.173E+03 1.243E+05 4.753E+02 2.698E+04 1.863E+02 1.270E+06 1.809E+02 1.656E+08
($\pm$1.705E+02) ($\pm$4.855E+04) ($\pm$5.596E+01) ($\pm$1.223E+04) ($\pm$9.315E-01) ($\pm$3.203E+05) ($\pm$1.733E+01) ($\pm$1.558E+07)
[…]
Method Sharp_Rid Different_Pow Schaffers_hig Composite_Gri Schwefel Gallagher_21P Katsuura Lunacek_bi_Ras
PSO 1.360E+03 1.928E+01 2.751E+01 5.665E+00 2.016E+01 6.659E+01 3.208E+00 3.461E+02
($\pm$1.372E+02) ($\pm$3.804E+00) ($\pm$4.309E+00) ($\pm$2.647E-01) ($\pm$1.457E+01) ($\pm$2.572E+00) ($\pm$2.201E-01) ($\pm$2.614E+01)
DE 7.012E+02 8.848E+00 1.536E+01 6.595E+00 2.439E+02 1.083E+01 3.000E+00 3.043E+02
($\pm$3.727E+01) ($\pm$7.387E-01) ($\pm$5.156E+00) ($\pm$1.962E-01) ($\pm$1.181E+02) ($\pm$7.941E-01) ($\pm$1.823E-01) ($\pm$3.679E+01)
CMAES 1.485E+03 2.591E+01 4.616E+01 4.878E+00 3.021E+00 7.824E+01 4.913E-01 2.853E+02
($\pm$1.474E+00) ($\pm$9.453E-02) ($\pm$4.374E-02) ($\pm$1.767E-01) ($\pm$6.518E-02) ($\pm$7.518E-02) ($\pm$1.826E-01) ($\pm$1.699E+00)
SAHLPSO 1.500E+03 1.707E+01 3.268E+01 8.649E+00 4.673E+03 4.442E+01 3.192E+00 4.744E+02
($\pm$3.082E+01) ($\pm$8.338E+00) ($\pm$4.166E+00) ($\pm$1.578E+00) ($\pm$1.368E+03) ($\pm$8.733E+00) ($\pm$7.748E-01) ($\pm$2.208E+01)
RNNOPT 2.314E+03 7.363E+01 7.426E+01 6.430E+00 2.606E+04 8.447E+01 7.372E+00 4.541E+02
($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00)
DEDQN 2.360E+03 4.823E+01 4.896E+01 2.184E+01 6.849E+04 7.642E+01 4.808E+00 8.033E+02
($\pm$1.659E+02) ($\pm$2.662E+00) ($\pm$5.639E+00) ($\pm$1.482E+00) ($\pm$1.161E+04) ($\pm$3.011E+00) ($\pm$3.796E-01) ($\pm$9.052E+01)
LES 2.895E+03 1.310E+02 1.341E+03 1.300E+03 7.007E+03 1.180E+03 1.303E+03 1.749E+03
($\pm$3.334E+01) ($\pm$1.310E+00) ($\pm$2.615E+00) ($\pm$4.597E-04) ($\pm$1.338E+03) ($\pm$1.277E+00) ($\pm$2.324E-01) ($\pm$1.423E+01)
LGA 2.821E+03 9.115E+01 6.749E+01 2.159E+01 9.618E+04 8.395E+01 3.231E+00 9.121E+02
($\pm$1.272E+02) ($\pm$2.051E+01) ($\pm$5.473E+00) ($\pm$3.009E+00) ($\pm$1.232E+04) ($\pm$1.719E+00) ($\pm$2.440E-01) ($\pm$1.066E+02)
GLHF 2.611E+03 6.510E+01 5.337E+01 2.541E+01 7.450E+04 8.305E+01 6.515E+00 8.946E+02
($\pm$1.152E+02) ($\pm$5.242E+00) ($\pm$5.011E+00) ($\pm$1.609E+00) ($\pm$4.114E+03) ($\pm$1.280E+00) ($\pm$1.767E-01) ($\pm$3.887E+01)
B2OPT 1.393E+03 1.921E+01 3.485E+01 2.156E+00 1.671E+03 7.715E+01 4.219E+00 3.267E+02
($\pm$1.978E+02) ($\pm$3.729E-01) ($\pm$1.291E+00) ($\pm$2.386E-01) ($\pm$8.537E+02) ($\pm$3.433E+00) ($\pm$2.024E-01) ($\pm$5.097E+00)
[…]
indicates ours variant enhanced with the proposed Mamba-based neural operator.
Compared Methods. We conduct a comparative analysis against representative algorithms to evaluate the effectiveness of the proposed L2E framework. This set encompasses both classical heuristics and learning-based optimizers. To situate our non-expansive unrolling paradigm within the broader L2O landscape, we incorporate two categories of learned algorithms. The first category consists of meta-learning-based neuro-evolution strategies, i.e., DEDQN [58], LES [28], and LGA [59]. The second category includes control-based solution generation models, i.e., RNNOPT [60], GLHF [61], and B2OPT [23]. Additionally, we employ established traditional algorithms, i.e., CMA-ES [11], DE [62], PSO [31], and SAHLPSO [63], which serve as performance anchors. We re-evaluate all competing methods within a unified experimental protocol and identical computational budgets, thereby ensuring a consistent assessment. Finally, we conduct all experiments on a single NVIDIA RTX A6000 GPU (PyTorch 2.0.1) and evaluate each task over 10 independent random seeds to report the mean and standard deviation.
V-B Performance on Numerical Benchmarks
We first examine the optimization capability of L2E on numerical landscapes where the meta-training and testing distributions align. This assessment covers two critical dimensions: precision on standard numerical problems (BBOB) and scalability to high dimensions (LSGO).
V-B1 Precision and Stability on Standard Benchmarks
Table I and Table II report the comparative results on BBOB-10D and BBOB-30D, respectively. L2E demonstrates a substantial advantage over state-of-the-art methods:
  • Performance Superiority and Architectural Efficacy: L2E establishes a consistent performance advantage. Specifically, it achieves the lowest mean error on 31 out of 32 test instances. On ill-conditioned and multimodal functions (e.g., Rosenbrock_rot, Lunacek_bi_Ras), L2E substantially outperforms strong baselines, such as GLHF and B2OPT. We attribute this escape capability to the non-expansive unrolling mechanism, which enforces a strict contractive constraint to prevent divergence in high-curvature valleys. Furthermore, the comparison between variants reveals that the Mamba-enhanced model (Ours) consistently improves upon the base model (Ours), particularly on non-separable landscapes (e.g., Ellipsoidal_hc). This result confirms that the structured state-space model offers a superior inductive bias for capturing complex variable couplings, in contrast to the generic attention or coordinate-wise mechanisms used in the standard version.
  • Convergence Trajectory Analysis: Figure 4 and Figure 5 illustrate the convergence behavior. L2E (red curves) exhibits a steeper initial descent and tighter confidence intervals. On unimodal functions (Bent Cigar), the model rapidly saturates to the optimum. This behavior demonstrates the effective exploitation of gradient-like information. Conversely, on multimodal tasks (Gallagher 21Peaks), L2E maintains steady progress without early stagnation. This balance confirms the efficacy of the Elite-Mask Fusion, which adaptively arbitrates between global exploration and local gradient refinement.
V-B2 Scalability to High-Dimensional Regimes
The results on the 1000D LSGO benchmark (Table III) further highlight the structural advantages of L2E. Specifically, L2E ranks first across all six representative high-dimensional functions. On the complex Shifted_Rastrigin, the model achieves a lower objective value, a result that surpasses LGA and GLHF. Existing Transformer-based optimizers (e.g., GLHF) encounter substantial bottlenecks in 1000D regimes due to the quadratic complexity of the self-attention mechanism. In contrast, our Mamba-based operator maintains linear complexity, which facilitates stable gradient propagation during meta-training. The ECDF curves in Fig. 6 and boxplots in Fig. 7 illustrate this distinction. While RNNOPT and DEDQN exhibit flatlining trajectories, which indicate gradient vanishing or saturation, L2E maintains a continuous descent. This behavior demonstrates the capability of L2E to navigate high-dimensional manifolds efficiently.
Attractive Sector
Bent Cigar
Buche Rastrigin
Composite Grie-Rosen
Different Powers
Discus
Ellipsoidal
Gallagher 21Peaks
Figure 4: Log-scaled convergence trajectories of our method and representative methods on the first 8 benchmark tasks from the BBOB-10D suite [54]. The x-axis denotes the number of function evaluations (FEs), while the y-axis reports the logarithm of the objective value (log cost), where lower values indicate superior performance. For clarity, only 8 representative tasks are visualized; Complete results on all 16 functions are provided in the supplementary material (i.e., Fig S1).
V-C Generalization to Real-World UAV Planning
A core contribution of L2E is its capability to generalize to unseen, real-world tasks without fine-tuning. We validate this Sim-to-Real transfer capability using the UAV Path Planning benchmark. Although the model is trained exclusively on synthetic BBOB-30D functions, L2E generalizes robustly to the 56 distinct UAV planning tasks (30D). As depicted in the radar plots of Fig. 8, L2E achieves the highest normalized performance across all seven terrain groups. The improvement margins over the second-best baseline range from 3.4% to 14.2%. This performance advantage stems from the proposed transfer mechanism. In contrast to reinforcement learning-based methods (e.g., DEDQN), which risk overfitting to specific landscape features encountered during training, L2E learns a generic descent operator that is governed by fixed-point properties. Consequently, the optimizer perceives the irregular constraints and noise profiles of the UAV simulator as intrinsic properties of the black-box function. This perception allows it to apply the learned contractive rules to navigate feasible regions.
V-D Neuroevolution for Robot Control
Finally, we assess the applicability of the framework to high-dimensional policy search tasks in the Neuroevolution (NE) benchmark [57]. This assessment represents a substantial domain shift from function optimization to control policy learning. The benchmark consists of six realistic robotic tasks across two domains, i.e., ant and humanoid. Specifically, each domain contains three complexity variants (ant-3, ant-4, ant-6; and humanoid-3, humanoid-4, humanoid-5). We benchmark the proposed method against GLEET [64], i.e., a specialized baseline that formulates neuroevolution within a reinforcement learning framework. GLEET trains an actor-critic agent to evolve population-level solutions via environmental interactions. We conduct this comparison to demonstrate that the learned optimizer can replace heuristic operators to optimize actor-critic evolutionary control policies.
Fig. 9 presents the comparative results, utilizing a normalized metric to enable intuitive visual comparison across tasks with distinct reward scales. The vertical axis represents deviations from a task-specific mean, rescaled by a common exponential factor: $y_{\text{plot}}=\frac{y-\text{mean}}{10^{\text{exp}}}.$ This transformation highlights the convergence dynamics around a central baseline, and it preserves cross-task comparability. On representative mid-difficulty tasks, such as ant-4 and humanoid-4, L2E exhibits steeper convergence trajectories and lower terminal costs than the baseline. Furthermore, the aggregated boxplots (Right) reveal that L2E achieves lower final objective values with consistently reduced variance across all six tasks. This result suggests that replacing heuristic population updates with the learned non-expansive operator stabilizes the policy learning process, thereby mitigating the instability typically associated with high-variance reward signals in RL environments.
Attractive Sector
Bent Cigar
Buche Rastrigin
Composite Grie-Rosen
Different Powers
Discus
Ellipsoidal
Gallagher 21Peaks
Figure 5: Log-scaled convergence trajectories of our method and representative methods on the first 8 benchmark tasks from the BBOB-30D suite [54]. The x-axis denotes the number of function evaluations (FEs), while the y-axis reports the logarithm of the objective value (log cost), where lower values indicate superior performance. For clarity, only 8 representative tasks are visualized; Complete results on all 16 functions are provided in the supplementary material (i.e., Fig S2).
TABLE III: High-Dimensional numerical evaluation (mean $\pm$ std) of representative algorithms on the LSGO-1000D. Each entry is calculated from 10 independent test runs. The best and second-best results are highlighted using bold and bold, respectively.
Method Shifted_Elliptic Shifted_Rastrigin Shifted_Ackley 7n1s_Shifted_Ell 7n1s_Shifted_Ras 7n1s_Shifted_Ack
PSO 3.689E+11 1.158E+05 2.164E+01 5.376E+13 7.802E+07 1.067E+06
($\pm$1.040E+10) ($\pm$2.302E+03) ($\pm$7.262E-03) ($\pm$1.431E+13) ($\pm$2.797E+06) ($\pm$1.919E+03)
DE 3.673E+11 1.236E+05 2.167E+01 3.722E+13 7.243E+07 1.071E+06
($\pm$5.347E+09) ($\pm$2.036E+03) ($\pm$4.901E-03) ($\pm$9.326E+12) ($\pm$2.896E+06) ($\pm$5.270E+02)
CMAES 2.098E+11 6.919E+04 2.162E+01 1.068E+14 5.580E+07 1.073E+06
($\pm$7.548E+05) ($\pm$1.941E+01) ($\pm$2.688E-03) ($\pm$3.869E+09) ($\pm$3.958E+04) ($\pm$2.152E+02)
SAHLPSO 1.877E+11 5.983E+04 2.162E+01 1.511E+13 3.510E+07 1.068E+06
($\pm$1.578E+09) ($\pm$1.677E+03) ($\pm$3.670E-02) ($\pm$9.400E+12) ($\pm$4.426E+06) ($\pm$1.274E+03)
RNNOPT 2.099E+11 4.764E+04 2.169E+01 1.111E+14 4.870E+07 1.076E+06
($\pm$0.000E+00) ($\pm$7.276E-12) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00)
DEDQN 3.585E+11 1.182E+05 2.166E+01 3.071E+13 6.578E+07 1.072E+06
($\pm$1.908E+10) ($\pm$3.626E+03) ($\pm$9.260E-03) ($\pm$1.101E+13) ($\pm$6.791E+06) ($\pm$1.731E+03)
LES 1.961E+11 4.707E+04 2.164E+01 2.778E+13 4.179E+07 1.069E+06
($\pm$1.948E+09) ($\pm$2.559E+02) ($\pm$7.111E-03) ($\pm$9.982E+12) ($\pm$1.164E+06) ($\pm$7.265E+02)
LGA 4.042E+11 1.262E+05 2.164E+01 5.189E+13 8.025E+07 1.068E+06
($\pm$2.897E+10) ($\pm$5.863E+03) ($\pm$5.471E-03) ($\pm$2.901E+13) ($\pm$9.417E+06) ($\pm$1.616E+03)
GLHF 3.645E+11 1.176E+05 2.166E+01 3.329E+13 6.923E+07 1.072E+06
($\pm$1.521E+10) ($\pm$3.311E+03) ($\pm$9.850E-03) ($\pm$1.083E+13) ($\pm$6.324E+06) ($\pm$1.332E+03)
B2OPT 3.564E+11 1.152E+05 2.166E+01 3.589E+13 7.028E+07 1.071E+06
($\pm$1.644E+10) ($\pm$4.849E+03) ($\pm$6.181E-03) ($\pm$1.583E+13) ($\pm$4.294E+06) ($\pm$2.221E+03)
[…]
VI Ablation Analysis and Mechanistic Insights
To disentangle the contribution of each algorithmic component and justify our design choices, we conduct a systematic ablation study on the BBOB-10D suite. This analysis is structured into two parts: (1) an investigation of the neural parameterization strategy for the evolution operator, specifically Shared versus Unshared Weights; and (2) a dissection of the combined efficacy of the core modules, i.e., ProxyGrad, SoftGate, and Mamba.
VI-A Evolutionary Curriculum via Unshared Parameterization
A pivotal architectural decision in deep unrolling is the strategy for sharing the meta-parameters $\mathcal{O}_{\mathrm{evo}}(\cdot;\bm{\omega})$ of the evolution operator $\mathcal{O}_{\mathrm{evo}}$ across time steps $k=1\dots K$.
  • Shared ($\bm{\omega}_{\text{shared}}$): This strategy ties a single set of weights across all unrolled blocks, which forces the operator to learn a time-invariant update rule.
  • Unshared ($\bm{\omega}^{k}$): This strategy allocates independent weights for each block, which allows the operator to evolve distinct behaviors at different stages of the optimization process.
Rot. Ackley
Rot. Rastrigin
Rot. Elliptic
Shifted Ackley
Figure 6: Empirical Cumulative Distribution Functions (ECDFs) of different methods on four representative LSGO-1000D [55]. The x-axis indicates the number of function evaluations required to reach a predefined target cost, while the y-axis reflects the cumulative proportion of runs that achieved this target. Each curve illustrates the probability that a given method solves a task within a certain cost budget. Complete results on all 6 functions are provided in the supplementary material (i.e., Fig S8).
Rot. Ackley
Rot. Rastrigin
Rot. Elliptic
Shifted Ackley
Figure 7: Boxplots of representative methods on LSGO-1000D [55]. Each box shows the distribution of results for one method, where the central line denotes the median, the box boundaries indicate the interquartile range, and whiskers represent the variability across trials. Complete results on all 6 functions are provided in the supplementary material (i.e., Fig S7).
Terrain 1–8
Terrain 9–16
Terrain 17–24
Terrain 25–32
Terrain 33–40
Terrain 41–48
Terrain 49–56
Figure 8: Normalized radar plots of representative methods on the UAV Path Planning Benchmark [56], which consists of 56 realistic planning tasks. The models are trained on BBOB-30D and directly tested in the real-world UAV scenario. The percentage value marked in each subplot denotes our method’s improvement margin over the second-best baseline.
Figure 9: Comparison results on the Neuroevolution for Robot Control benchmark [57], comprising six high-dimensional robot control tasks from two domains: ant (ant-3, ant-4, ant-6) and humanoid (humanoid-3, humanoid-4, humanoid-5). The left and center plots depict convergence curves on two representative mid-difficulty tasks, where the $y$-axis represents normalized deviations from a task-specific mean, scaled by a shared exponential factor (e.g., $(\Delta\text{cost from }1.15096\times 10^{3}$, scaled by $10^{-3})$). The right plot shows boxplots of final objective values (log scale) aggregated across all six tasks.
VI-A1 Statistical Superiority of Unshared Design
The unshared parameterization strategy consistently outperforms the Shared baseline across all metrics. Table IV presents the aggregate performance across varying evaluation budgets. This advantage amplifies as the trajectory length increases. Specifically, at $\mathrm{MaxFEs}=20{,}000$ ($\mathrm{EMS}=200$), the Unshared model reduces the mean error by approximately 96% compared to the Shared version. This empirical gap suggests that a static operator is insufficient to handle the dynamic requirements of a complete optimization lifecycle. Conversely, the allocation of unique $\bm{\omega}$ to each block enables stage-wise specialization of the update dynamics. This capability is increasingly beneficial as $\mathrm{EMS}$ grows and the learning trajectory lengthens. Consequently, our final model adopts the separate per-block meta-parameters design for $\mathcal{O}_{\mathrm{evo}}$. Although this incurs a marginally higher training cost, it yields superior accuracy and stability under practical evaluation budgets.
TABLE IV: Ablation on Evolution Neural Block weight sharing on BBOB-10D. Each side aggregates over 16 functions using only the Obj metric (lower is better). Ranks are computed per function across the 8 configurations with average-tie handling. Std is computed across the 16 functions per configuration using population statistics. EMS denotes the number of epochs/main steps, computed as $\big\lceil\mathrm{MaxFEs}/\mathrm{NP}\big\rceil$ with $\mathrm{NP}=100$ (i.e., ems = (MaxFEs + NP ) // NP - 1). Best overall is bold.
Shared = True: $\mathcal{O}_{\mathrm{evo}}$ uses shared meta-parameters $\bm{\omega}$ Shared = False: $\mathcal{O}_{\mathrm{evo}}$ uses separate per-block meta-parameters $\bm{\omega}$
MaxFEs EMS Mean $\downarrow$ Median $\downarrow$ GeoMean $\downarrow$ Std Dev $\downarrow$ Avg. Rank $\downarrow$ MaxFEs EMS Mean $\downarrow$ Median $\downarrow$ GeoMean $\downarrow$ Std Dev $\downarrow$ Avg. Rank $\downarrow$
2,000 20 2.131E+06 1.375E+02 4.281E+02 8.162E+06 8.00 2,000 20 5.442E+05 3.706E+01 1.106E+02 2.088E+06 5.06
5,000 50 5.294E+05 4.777E+01 1.226E+02 2.000E+06 5.62 5,000 50 2.800E+05 3.091E+01 7.729E+01 1.070E+06 3.31
10,000 100 8.150E+05 4.469E+01 1.088E+02 3.126E+06 5.00 10,000 100 1.350E+05 2.768E+01 5.521E+01 5.139E+05 2.00
20,000 200 9.302E+05 4.850E+01 1.169E+02 3.581E+06 6.00 20,000 200 3.506E+04 2.239E+01 3.879E+01 1.312E+05 1.00
Figure 10: Convergence behaviors under the shared vs. unshared parameterization of the neural evolution operator $\mathcal{O}_{\mathrm{evo}}$. Left two: log-cost convergence on the Ellip-HC problem; Right two: aggregated normalized cost curves over all 16 BBOB-like test problems. Additional convergence curves for all problems are provided in the supplementary material (i.e., Figs. S9 and S10).
VI-A2 Stage-wise Specialization
Visual analysis confirms that the unshared parameterization enables an implicit evolutionary curriculum. Fig. 10 provides a visual explanation of this phenomenon. In the early stage, i.e., exploration, both configurations descend rapidly. However, in the late stage, i.e., exploitation, the Shared model (brown curves) flattens out prematurely. It functions as a stationary operator, which prevents mode switching. In contrast, the unshared model (red/purple curves) continues to refine the solution. The unshared parameterization effectively learns an implicit evolutionary curriculum. Specifically, early blocks specialize in global exploration to escape local basins, whereas later blocks shift towards precise local refinement. This time-dependent specialization is pivotal for high-precision convergence in black-box optimization.
VI-B Deconstructing Component Synergy
We isolate the contributions of the three core architectural components of L2E: (i) ProxyGrad, which provides gradient injection; (ii) SoftGate, which enables adaptive fusion; and (iii) Mamba, which functions as the structured operator. We compare the full model, denoted as Ours, against variants where we individually remove or replace each component.
VI-B1 Quantitative Impact Analysis
Table V quantifies the performance degradation that results from the removal of individual components.
  • Impact of Gradient Guidance (w/o ProxyGrad): The removal of proxy gradients results in the most substantial performance deterioration, as the error increases by $\sim 103\times$. This outcome confirms that purely evolutionary search is sample-inefficient in high-dimensional spaces. Consequently, proxy gradients provide the descent momentum that is necessary to guide the population.
  • Impact of Adaptive Fusion (w/o SoftGate): The disablement of the gating mechanism, i.e., the enforcement of a deterministic combination, degrades performance by $\sim 161\times$. This result demonstrates the necessity of selective utilization. Specifically, the solver must dynamically decide when to trust the gradient for exploitation and when to rely on evolutionary proposals for exploration.
  • Impact of Architecture (w/o Mamba): The replacement of Mamba with a standard MLP/Attention block increases error by $\sim 6.8\times$. Although this degradation is less severe than the removal of gradients, it confirms that Mamba enables the modeling of long-range dependencies and coupled variables. This capacity is essential for breaking the precision ceiling on complex landscapes.
VI-B2 Convergence Dynamics
Fig. 11 and Fig. 12 visualize the per-problem behavior and convergence stability. On ill-conditioned functions, such as Ellipsoidal_hc, the Full model (red) achieves the lowest terminal cost and the tightest variance (Fig. 12, Right). In contrast, the w/o SoftGate variant exhibits high variance, which indicates that rigid updates lead to instability.
The proposed components function in concert. Specifically, ProxyGrad provides speed, Mamba ensures structural adaptation to landscape geometry, and SoftGate serves as a robust arbiter. The removal of any single component undermines this synergy, thereby leading to slower convergence or stagnation. Thus, L2E represents a cohesive system rather than a simple sum of parts. Its superior performance stems from the stage-wise curriculum, which is enabled by unshared unrolling, and the adaptive interplay between gradient-based and evolutionary search. This interplay is structurally supported by the non-expansive Mamba operator.
TABLE V: Overall objective on 16 BBOB problems. The first three columns indicate whether each component is enabled (✓) or disabled (✗). In the Mean ($\downarrow$) column, the parentheses report the reduction of Ours (Full), Item 4 relative to each ablated variant, computed as $(x-\text{Full})/x$. Best is marked with bold.
Item ProxyGrad SoftGate Mamba Mean ($\downarrow$) Std Dev
1 5.294E+05  ($\downarrow$99.04%) 2.007E+06
2 8.213E+05  ($\downarrow$99.38%) 3.150E+06
3 3.506E+04   ($\downarrow$85.46%) 1.312E+05
4 5.098E+03(ref.) 1.845E+04
Figure 11: Log-scale objective values (lower is better) on the first eight BBOB problems under four ablation configurations.
Figure 12: Ablation on convergence and terminal cost across four variants (w/o ProxyGrad, w/o SoftGate, w/o Mamba, and Ours (Full)). Left: log objective gap vs. function evaluations (FEs) on two representative problems. Right: box-whisker statistics of terminal cost across runs.
VII Conclusion
In this paper, we proposed L2E, a unified framework that approaches evolutionary optimization via neural unrolling. L2E enforces strict fixed-point contraction within a bilevel meta-learning scheme. This architecture seamlessly integrates the expressiveness of deep learning with the convergence guarantees of classical optimization. Empirical results across high-dimensional benchmarks and robot control tasks demonstrate the scalability and generalization capabilities. Limitations: However, the reliance on structured neural operators introduces computational overheads. Consequently, this computational cost restricts applicability in resource-constrained or low-latency environments. Future work will focus on developing lightweight neural operators that are compatible with real-time execution. Additionally, we aim to explore sample-efficient first-order algorithms to enhance inference speed. Finally, we plan to extend this non-expansive unrolling paradigm to multi-objective and constrained optimization scenarios.

### v2

@@
-Learning to Evolve with Convergence Guarantee via Neural Unrolling
+Learning to Evolve for Optimization via Stability-Inducing Neural Unrolling
 Abstract
-The transition from hand-crafted heuristics to data-driven evolutionary algorithms faces a fundamental dilemma: achieving neural plasticity without sacrificing mathematical stability.
-Emerging learned optimizers demonstrate high adaptability.
-However, they often lack rigorous convergence guarantees.
-This deficiency results in unpredictable behaviors on unseen landscapes.
-To address this challenge, we introduce Learning to Evolve (L2E), a unified bilevel meta-optimization framework.
-This method reformulates evolutionary search as a Neural Unrolling process grounded in Krasnosel’skii-Mann (KM) fixed-point theory.
-First, L2E models a coupled dynamic system in which the inner loop enforces a strict contractive trajectory via a structured Mamba-based neural operator.
-Second, the outer loop optimizes meta-parameters to align the fixed point of the operator with the target objective minimizers.
-Third, we design a gradient-derived composite solver that adaptively fuses learned evolutionary proposals with proxy gradient steps, thereby harmonizing global exploration with local refinement.
-Crucially, this formulation provides the learned optimizer with provable convergence guarantees.
-Extensive experiments demonstrate the scalability of L2E in high-dimensional spaces and its robust zero-shot generalization across synthetic and real-world control tasks.
-These results confirm that the framework learns a generic optimization manifold that extends beyond specific training distributions.
+Evolutionary algorithms serve as a powerful paradigm for tackling optimization challenges, yet their reliance on manually engineered heuristics inherently limits their adaptability across diverse landscapes.
+However, the transition from the hand-crafted heuristics to data-driven algorithms faces a fundamental dilemma: achieving neural plasticity without sacrificing algorithmic stability.
+Although learned optimizers offer high adaptivity, their unconstrained update rules often result in unstable dynamics and brittle generalization on unseen landscapes.
+To address this challenge, this paper proposes Learning to Evolve (L2E), a bilevel meta-optimization framework that learns evolutionary search via stability-inducing neural unrolling.
+First, L2E reformulates population evolution as an unrolled fixed-point iteration via a structured neural operator.
+In this design, the inner loop imposes a stability-biased update structure, while the outer loop meta-trains the operator to produce effective search trajectories across tasks.
+Second, to balance global exploration with local refinement, a gradient-derived composite solver adaptively fuses learned evolutionary proposals with proxy numerical guidance in a differentiable manner.
+Extensive experiments on synthetic benchmarks and real-world control tasks demonstrate that L2E achieves substantial optimization performance, scales to high-dimensional problems, and exhibits robust zero-shot transfer across diverse test distributions.
 I Introduction
-Evolutionary Algorithms (EAs) constitute a core paradigm in computational intelligence for solving complex black-box optimization problems [1, 2, 3, 4, 5].
-These algorithms evolve a population of candidate solutions through selection, recombination, and perturbation.
-This mechanism enables them to explore highly nonconvex, multimodal, and noisy landscapes without relying on explicit gradient information.
-Consequently, this paradigm plays a foundational role in computational science and engineering.
-It underpins diverse applications such as hyperparameter tuning in machine learning [6], aerodynamic shape design [7], molecular conformation prediction [8], and the control of physical systems [9].
-Specifically, solvers in these scenarios must navigate high-dimensional and noisy search spaces under tight evaluation budgets.
-They must simultaneously achieve effective exploration and reliable convergence to high-quality solutions [10].
-Optimization research has established two principal families over the past decades: gradient-free, population-based Evolutionary Algorithms (EAs) and model-based Bayesian Optimization (BO).
-The former family, which includes Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [11], Differential Evolution (DE) [12], and various swarm-based methods [13], is noted for conceptual simplicity and global exploration.
-Conversely, the latter family is widely adopted for sample efficiency in computationally expensive contexts [14].
-However, both families rely on hand-crafted operators and heuristics.
-Furthermore, they lack a unified theoretical foundation for convergence and stability.
-This inherent limitation constrains their efficiency in high-dimensional spaces.
-Consequently, this results in a fundamental lack of adaptability to increasingly complex modern workloads.
-A natural response to these limitations is the shift from manually designing operators to data-driven and learnable evolutionary optimizers [15, 16, 17].
-Within the Learning to Optimize (L2O) paradigm, recent studies leverage Recurrent Neural Networks (RNNs) and Transformers as controllers, or employ set-based architectures to model optimization trajectories [18].
-In parallel, meta-evolution approaches treat mutation-crossover-selection policies as learnable objects.
-These methods search over expressive families of attention-based, diffusion-inspired, or transformer-style operators [19, 20, 21].
-Specifically, they seek to automatically discover evolutionary strategies by learning from a task distribution, such as LQD [22] and the Evolution Transformer [23].
-However, a critical gap remains in this transition: the emerging learned optimizers typically face a fundamental dilemma between plasticity and stability.
-Most existing approaches treat the optimizer as a black-box policy network trained via reinforcement learning to simply minimize trajectory loss.
-While they demonstrate exceptional plasticity on training tasks, they frequently degenerate into mathematically unconstrained processes.
-Lacking intrinsic geometric constraints, they are prone to overfitting the training landscape topology, figuratively “memorizing the map” rather than “learning the compass”.
-Consequently, when deployed on unseen or out-of-distribution tasks, these solvers often exhibit wandering trajectories, catastrophic divergence, or stagnation, failing to provide the stability and rigorous convergence assurances characterizing classical EAs.
-Consequently, this limitation motivates us to rethink the fundamental design of learned optimizers: how can we engineer a learned evolutionary optimizer that possesses the adaptive intelligence of neural networks without sacrificing the rigorous convergence laws of classical control?
-To address this challenge, we propose to ground the design of learned EAs in fixed-point theory.
-We argue that the fragility of current learned optimizers stems from unconstrained operator learning.
-A robust evolutionary operator should not merely predict the next step based on heuristics; rather, it must mathematically constitute a contraction mapping.
-The repeated application of this mapping drives the population towards equilibrium.
-Motivated by this insight, we introduce Learning to Evolve (L2E), a unified framework that reformulates evolutionary search as a neural unrolling process.
-Unlike approaches that view the optimizer merely as a meta-learner, L2E models the optimizer as a learnable dynamical system governed by the classical Krasnosel’skiĭ-Mann (KM) iteration [24].
-The core philosophy is to enforce a physics of descent within the neural architecture.
-In the inner loop, the population update serves as an $\alpha$-averaged operator composed of a learned component and a numerical proxy.
-This structure enforces non-expansiveness by design.
-Consequently, the evolution trajectory strictly contracts towards a fixed point, ensuring stability regardless of the landscape’s complexity.
-In the outer loop, the meta-learner aligns the operator’s fixed point with the target objective minimizers.
-This alignment effectively shapes the attraction basin of the dynamics.
-To implement this efficiently in high-dimensional spaces, we propose a structured Mamba-based operator, where spectral normalization constraints are imposed.
-By leveraging the linear complexity of State Space Models (SSMs), this design overcomes the quadratic bottleneck of Transformers in population-wise interaction complexity, capturing long-range dependencies dynamics efficiently.
-This architectural formulation yields the behavioral contrast illustrated in Fig.
+Evolutionary Algorithms (EAs) are a core paradigm in computational intelligence for solving complex black-box optimization problems [67, 46, 65, 8, 5].
+By evolving a population through selection, recombination, and perturbation, EAs can explore highly nonconvex, multimodal, and noisy landscapes without relying on explicit gradient information.
+As a result, they underpin a wide range of applications, including hyperparameter tuning [14], aerodynamic shape design [24], molecular conformation prediction [51], and the control of physical systems [15].
+In these scenarios, solvers must operate under tight evaluation budgets while navigating high-dimensional and noisy search spaces, requiring both effective exploration and reliable progress toward high-quality solutions [9].
+Over the past decades, black-box optimization has been dominated by two major families: population-based evolutionary search and model-based Bayesian optimization (BO).
+The former emphasizes exploration through handcrafted operators and heuristics, whereas the latter focuses on sample efficiency via surrogate modeling in expensive evaluation settings [12].
+Despite their empirical success, a common bottleneck is that the update dynamics are largely driven by manually designed heuristics.
+Consequently, their behavior is difficult to characterize and control in a unified way across varying scales, conditioning, and dimensionalities.
+This issue becomes particularly pronounced in high-dimensional regimes and when the optimizer is deployed on unseen tasks, where stable progress under limited evaluations is challenging.
+A natural response is the transition from manually designing operators to data-driven and learnable evolutionary optimizers [62, 70, 66, 63].
+Recent studies in learning to optimize (or meta-evolution) learn mutation-crossover-selection policies from a task distribution.
+Specifically, these methods employ neural controllers (e.g., RNNs, Transformers) or set-based architectures to model population updates [6, 3, 44, 45, 42].
+Representative examples include learning-driven operator discovery and attention-based evolutionary strategies, such as LQD [4] and the evolution transformer [32].
+However, a critical gap remains: learned optimizers often struggle to reconcile plasticity with stability in a way that transfers reliably across tasks.
+Most existing approaches learn update rules as unconstrained policies that minimize trajectory loss on training problems.
+While such policies can be highly effective in-distribution, they may induce unstable dynamics on unseen landscapes, such as oscillation, premature stagnation, or sensitivity to nuisance shifts in scale and conditioning.
+In other words, the learned update rule may implicitly overfit the training landscape topology, i.e., “memorizing the map” rather than “learning the compass”.
+This tendency results in brittle generalization under distribution shifts.
+This observation motivates a different design principle: rather than learn an unconstrained update policy, the optimizer should learn an iterative dynamical system with an explicit stability-inducing bias.
+In response, we propose Learning to Evolve (L2E), a bilevel meta-optimization framework that learns evolutionary search through stability-inducing neural unrolling.
+Concretely, L2E models population evolution as an unrolled iterative solver [17, 49].
+The inner loop follows a stability-biased update structure (e.g., averaged/residual updates with bounded transformations), while the outer loop meta-trains the operator to produce effective search trajectories.
+To balance global exploration with local refinement, L2E further introduces a gradient-derived composite solver.
+This solver adaptively fuses learned evolutionary proposals with proxy numerical guidance in a differentiable manner, enabling controlled progress under limited evaluations.
+From an architectural perspective, L2E parameterizes the learned evolutionary operator with a structured Mamba-based design to model population-wise interactions.
+By combining state-space modeling with lightweight routing over population statistics, the operator provides a scalable inductive bias for meta-learned evolutionary search.
+This stability-inducing formulation yields the behavioral contrast illustrated in Fig.
 1.
-While conventional learned optimizers often succumb to overfitting the training landscape, which results in unpredictable or unstable trajectories on unseen tasks, L2E induces a stable contraction field where the learned fixed point is explicitly aligned with the target optimum.
-Consequently, the proposed L2E avoids the pitfall of memorizing the map by learning a generalized descent rule, thereby offering a robust solution for general-purpose adaptation.
+Specifically, while unconstrained learned heuristics may overfit training landscapes and exhibit unstable trajectories on unseen problems, L2E promotes controlled search dynamics by learning structured evolutionary updates via neural unrolling.
+This framework explicitly emphasizes stable dynamics, scalability to high-dimensional settings, and strong empirical generalization under distribution shifts.
 Our main contributions are:
-  • Theory-Driven Formulation: We introduce the Generalized Neural Unrolling (GNU) paradigm.
-This framework redefines evolution as a curriculum-driven unrolling process under bilevel fixed-point constraints.
-Furthermore, we establish a joint convergence guarantee to bridge neural plasticity with theoretical optimization rigor.
-  • Efficient Composite Solver: We design a Gradient-derived Evolutionary Composite (GEC) solver, which utilizes a Mamba architecture and an elite-mask fusion mechanism.
-This design addresses the scalability limitations of attention-based optimizers.
-Consequently, it enables stable meta-evolution in high-dimensional regimes where traditional L2O methods struggle.
-  • Rigorous Theoretical Analysis: We provide a fixed-point theoretic analysis establishing the finite-time convergence rate, asymptotic consistency, and bilevel optimality.
-These guarantees ensure the learned optimizer follows a predictable trajectory, avoiding the instability inherent in unconstrained meta-learning.
+  • Stability-Inducing Unrolling Formulation: We propose Learning to Evolve (L2E), a bilevel meta-optimization framework that learns evolutionary search via stability-inducing neural unrolling.
+The inner loop enforces structured population updates that are stability-oriented, while the outer loop meta-trains the operator to produce effective search trajectories across tasks.
+  • Structured Operator and Composite Solver: We develop a Gradient-derived Evolutionary Composite (GEC) solver that adaptively fuses learned evolutionary proposals with proxy numerical guidance in a differentiable manner.
+We further introduce a structured Mamba operator to capture population-wise interactions, providing a scalable inductive bias in high-dimensional regimes.
+  • Analysis and Empirical Validation: We analyze how the proposed unrolling structure induces controlled optimization dynamics, and validate L2E on synthetic benchmarks and real-world control tasks.
+Results show strong performance, scalability to high dimensions, and robust zero-shot generalization under distribution shifts.
 The remainder of the paper is organized as follows.
@@
 Figure 1: Conceptual contrast.
-Left: Naive optimizers memorize the map, leading to unstable wandering on unseen landscapes.
-Right: L2E enforces a physics of descent via non-expansive unrolling, inducing a stable contraction field where the learned fixed-point is aligned with the target optimum.
+Left: Unconstrained learned heuristics may overfit training landscapes, resulting in unstable trajectories on unseen problems.
+Right: L2E learns a stability-inducing unrolled evolutionary dynamics with bilevel meta-training, shaping the search trajectory toward controlled progress and improved robustness under distribution shifts.
 II Related Work
 II-A Learning to Optimize
-Learning to Optimize (L2O) blends the prescriptive rigor of classical optimization with the inductive power of machine learning [18].
+Learning to Optimize (L2O) combines the rigor of classical optimization with the capabilities of machine learning [3].
 Specifically, rather than relying on hand-crafted heuristics, L2O seeks to parameterize update rules directly from data.
@@
 This methodology has further evolved into meta-learning frameworks that train optimizers to adapt across diverse tasks [45, 27].
-Notable implementations include reinforcement learning controllers for scheduling [19] and gradient-guided mutation operators [23].
+Related directions include reinforcement learning controllers for scheduling [44] and gradient-guided mutation operators [32].
 Despite these advances, existing L2O approaches primarily function as unconstrained policy networks.
 Specifically, they lack intrinsic geometric constraints.
-This absence implies that there is no theoretical assurance of convergence or stability.
-Consequently, this deficiency frequently leads to unpredictable behavior when the model transfers to out-of-distribution landscapes, thereby highlighting the necessity for mathematically grounded architectures that enforce stability by design.
+This deficiency often leads to unstable optimization dynamics and brittle generalization under distribution shifts.
+Consequently, these limitations highlight the need for structure-aware learned optimizers that promote controlled update behaviors.
 II-B Learnable Evolutionary Algorithms
@@
 However, their reliance on hand-crafted, non-differentiable rules limits their flexibility in high-dimensional or ill-conditioned landscapes.
-To enhance adaptability, recent research has progressively integrated learnability into evolutionary frameworks [32, 33, 34].
+To enhance adaptability, recent research has progressively integrated learnability into evolutionary frameworks [71, 60, 29, 61, 28].
 Initial efforts utilized surrogates, such as Gaussian Processes, to approximate objective landscapes [40].
@@
 Despite this progress, learnable EAs face substantial challenges.
-Most existing works optimize isolated components rather than the unified evolutionary loop.
-This isolated optimization results in a fragmented learning process.
-Furthermore, current methods often lack a consistent mathematical framework to integrate gradient propagation with population search.
-Consequently, achieving robust zero-shot generalization and scalability in high-dimensional regimes remains a persistent difficulty.
+Most methods optimize isolated components, rather than the unified evolutionary loop, leading to fragmented training and limited end-to-end coordination between gradient guidance and population search.
+As a result, robust zero-shot generalization and scalability to high-dimensional regimes remain difficult to achieve.
 II-C Neural Unrolling
@@
 Furthermore, in the realm of AutoML, differentiable approaches for Neural Architecture Search (NAS) implicitly leverage the unrolling concept to optimize structural parameters via bilevel gradients [34, 35, 49].
-Consequently, it synergizes the interpretability of model-based methods with the adaptability of deep networks, thereby achieving robust performance across diverse inverse problems.
-In contrast to its prevalence in signal processing, NU remains underutilized within evolutionary computation.
-While differentiable EAs exist, they typically rely on ad hoc relaxations rather than systematic unrollings of fixed-point iterations.
-Thus, the field lacks a framework capable of unifying bilevel meta-learning with rigorous population dynamics guarantees.
+Consequently, it synergizes the interpretability of model-based methods with the adaptability of deep networks, thereby achieving robust performance across diverse problems.
+Despite its prevalence in signal processing, NU remains underutilized within evolutionary computation.
+Although differentiable EAs exist, they typically rely on ad hoc relaxations rather than systematic unrollings of fixed-point iterations.
+Consequently, the field lacks a framework that unifies bilevel meta-learning with stability-oriented population dynamics in a principled and trainable manner.
 II-D Discussion
-Existing methodologies navigate a critical trade-off between stability and adaptability.
-While classical EAs provide theoretical stability, they often lack the adaptability requisite for navigating complex landscapes efficiently.
-Conversely, emerging L2O approaches offer data-driven flexibility; however, they frequently exhibit instability and poor generalization to out-of-distribution landscapes, primarily due to the absence of intrinsic geometric constraints.
+Existing methodologies navigate a trade-off between stability and adaptability.
+While classical EAs provide theoretical stability, they often lack the adaptability for navigating complex landscapes efficiently.
+Conversely, emerging L2O methods offer data-driven flexibility; however, they frequently exhibit instability and poor generalization to out-of-distribution landscapes, primarily due to the absence of intrinsic geometric constraints.
 To reconcile these diverging paradigms, we propose the L2E that reformulates the evolutionary loop as a neural unrolling process.
-This formulation is expected to synergize the expressivity of deep networks with the rigorous convergence properties of classical optimization.
-Figure 2: Overall framework of the proposed L2E paradigm.
-Left: Visualizes the interaction where the outer-loop meta-learner reshapes the landscape to align the inner-loop fixed point with the target optimum.
-Middle: Illustrates the solving process, where the evolutionary search is unrolled into a differentiable pipeline to learn stable trajectories.
-Right: Details the non-expansive unrolling operator integrates evolutionary and numerical solvers to form a differentiable evolution dynamics.
+This formulation integrates the expressivity of deep networks with controlled optimization dynamics.
+Consequently, it improves robustness under distribution shifts.
+Figure 2: Overall framework of L2E.
+Left: L2E is formulated as a bilevel meta-optimization problem where the outer loop updates meta-parameters to shape the induced evolution dynamics.
+Middle: The evolutionary search is unrolled into a differentiable pipeline via AutoDiff to learn stable and effective search trajectories.
+Right: L2E implements a stability-inducing operator design that combines a learned evolutionary operator with numerical guidance, promoting controlled population updates.
 III Methodology
-In this section, we present L2E, a framework designed to unify heuristic search and mathematical optimization.
-The overall architecture functions as a bilevel dynamical system, as visualized in Fig.
+The overall architecture operates as a bilevel dynamical system, as visualized in Fig.
 2.
-To instantiate this paradigm, we first introduce the Generalized Neural Unrolling (GNU) formulation, which provides the theoretical backbone.
-Second, we detail the Gradient-derived Evolutionary Composite (GEC) solver, which harmonizes gradient-based exploitation with population-based exploration.
-Finally, we propose a Structured Mamba Neural Operator to efficiently parameterize this process in high-dimensional spaces, thereby ensuring both computational scalability and theoretical stability.
+To instantiate this paradigm, the methodology comprises three core components.
+First, the Generalized Neural Unrolling (GNU) formulation establishes the theoretical foundation.
+Second, the Gradient-derived Evolutionary Composite (GEC) solver integrates gradient-based exploitation with population-based exploration.
+Finally, the structured Mamba neural operator efficiently parameterizes the optimization process in high-dimensional spaces, thereby improving computational scalability and promoting controlled optimization dynamics.
 III-A Generalized Neural Unrolling Paradigm
-Conventional black-box optimizers often rely on heuristic mechanisms that lack theoretical guarantees.
-To address this limitation, we reinterpret the optimization process through the lens of non-expansive operator theory.
-Specifically, we utilize the KM fixed-point formulation [24].
+Conventional black-box optimizers often rely on heuristic mechanisms that can yield unstable dynamics on unseen landscapes.
+To promote controlled update behaviors, we reinterpret population evolution as a structured operator iteration, adopting a fixed-point inspired unrolling view based on the KM formulation [64].
 Unlike methods that directly search for optima, the proposed approach optimizes the parameters of a learnable fixed-point operator.
@@
 Formally, let $f:\mathcal{X}\to\mathbb{R}$ denote the black-box objective function.
-We introduce a learnable non-expansive operator $\mathcal{NU}(\cdot;\bm{\omega})$, parameterized by $\bm{\omega}\in\bm{\Omega}$, whose fixed point approximates the minimizer of $f$.
+We introduce a learnable non-expansive-inspired operator $\mathcal{NU}(\cdot;\bm{\omega})$, parameterized by $\bm{\omega}\in\bm{\Omega}$, whose fixed point approximates the minimizer of $f$.
 Consequently, we formulate this framework as a constrained bilevel optimization problem [36]:
 $\min_{\mathbf{x}\in\mathcal{X},\bm{\omega}\in\Omega}\ell_{\mathtt{meta}}(\mathbf{x};\bm{\omega},\mathtt{Opt}),s.t.\;\mathbf{x}\in\operatorname{Fix}(\mathcal{NU}(\cdot;\bm{\omega})),$ (1)
-where $\ell_{\mathtt{meta}}(\mathbf{x};\bm{\omega},\mathtt{Opt})$ denotes a meta-level surrogate objective that encodes task rewards, diversity, or other downstream fitness metrics.
+where $\ell_{\mathtt{meta}}(\mathbf{x};\bm{\omega},\mathtt{Opt})$ is the meta-objective that evaluates the population state $\mathbf{x}$ induced by the optimizer parameter $\bm{\omega}$ under a prescribed evaluation protocol $\mathtt{Opt}$.
 Furthermore, $\text{Fix}(\mathcal{NU}):=\{\bm{\mathbf{x}}\mid\bm{\mathbf{x}}=\mathcal{NU}(\bm{\mathbf{x}};\bm{\omega})\}$ defines the set of fixed points.
-In this formulation, $\bm{\omega}$ represents the outer-level parameters defining the optimization algorithm, whereas $\bm{\mathbf{x}}$ denotes the inner-level population of candidate solutions.
-To ensure contractive behavior, we construct $\mathcal{NU}$ as an $\alpha$-averaged operator.
-Specifically, this formulation employs a convex combination of the identity and a learnable mapping $\mathcal{O}(\cdot,\bm{\omega})$, thereby yielding a KM-style operator:
+Here, $\bm{\omega}$ are outer-level meta-parameters that parameterize the optimizer, whereas $\mathbf{x}$ denotes the inner-level population state of candidate solutions.
+To introduce a stability-biased update structure, we construct $\mathcal{NU}$ as an $\alpha$-averaged (KM-style) operator.
+Specifically, this formulation employs a convex combination of the identity and a learnable composite mapping $\mathcal{O}(\cdot,\bm{\omega})$, i.e.,
 $\bm{\mathbf{x}}^{k+1}=\mathcal{NU}(\mathbf{x}^{k};\bm{\omega})=(1-\alpha)\cdot\mathbf{x}^{k}+\alpha\cdot\mathcal{O}(\mathbf{x}^{k};\bm{\omega}),$ (2)
@@
 The operator $\mathcal{NU}(\cdot;\bm{\omega})$ is defined as a composition of two components, modulated by an update relaxation parameter $\alpha\in(0,1]$.
-Specifically, $\mathcal{O}_{\text{num}}$ denotes a numerical solver, such as smoothing or projection, while $\mathcal{O}_{\text{evo}}$ encapsulates a learnable search policy.
-By enforcing non-expansiveness on $\mathcal{O}$, the GNU paradigm ensures that the evolution trajectory $\mathbf{x}^{0}\to\dots\to\mathbf{x}^{K}$ follows a strictly contractive path.
-Consequently, this formulation provides a theoretical safeguard against divergence, which addresses a common failure mode in standard learned optimizers.
+$\mathcal{O}_{\mathrm{evo}}$ is a learnable evolutionary proposal operator, and $\mathcal{O}_{\mathrm{num}}$ provides proxy numerical guidance (e.g., projection or smoothing) to regulate the update behavior.
+Together, the averaging step and the composite structure impose a stability-inducing bias on the unrolled trajectory, which empirically reduces the brittle dynamics often observed in unconstrained learned optimizers.
 This formulation gives rise to a semantically structured learning process: i) Inner-Level (Evolve): For fixed algorithm parameters $\bm{\mathbf{\omega}}$, the inner loop performs K-step unrolled evolution of an initial population $\bm{\mathbf{x}}^{0}$.
@@
 6  for $k=1$ to $K$ do
-     /* Step 1: Non-expansive Neural Evolution (KM Iteration) */
-7     $\mathbf{d}_{IL}^{k}\leftarrow\mathbf{x}^{k-1}+\alpha\cdot\left(\mathcal{O}_{\mathrm{evo}}(\mathbf{x}^{k-1},\bm{\omega})-\mathbf{x}^{k-1}\right)$
+     /* Step 1: Stability-inducing Neural Evolution (KM Iteration) */
+7     $\mathbf{d}_{IL}^{k}\leftarrow(1-\alpha)\mathbf{x}^{k-1}+\alpha\cdot\mathcal{O}_{\mathrm{evo}}(\mathbf{x}^{k-1},\bm{\omega})$
 8
@@
 14
-    /* Step 3: Adaptive Manifold Rectification */
-     $\Delta f^{k}\leftarrow f(\mathbf{d}_{OL}^{k})-f(\mathbf{d}_{IL}^{k})$;    $\bm{M}^{k}\leftarrow\sigma(-\Delta f^{k}/\tau)$
-     // Soft Gating
-15     $\mathbf{x}^{k}\leftarrow\text{Proj}_{\mathcal{X},\bm{P}_{\bm{\omega}}}\left(\bm{M}^{k}\odot\bm{d}_{OL}^{k}+(1-\bm{M}^{k})\odot\bm{d}_{IL}^{k}\right)$
+    /* Step 3: Soft-gated Fusion */
+15     $\Delta f^{k}\leftarrow f(\mathbf{d}_{OL}^{k})-f(\mathbf{d}_{IL}^{k})$;    $\bm{M}^{k}\leftarrow\sigma(-\Delta f^{k}/\tau)$ $\mathbf{x}^{k}\leftarrow\text{Proj}_{\mathcal{X},\bm{P}_{\bm{\omega}}}\left(\bm{M}^{k}\odot\bm{d}_{OL}^{k}+(1-\bm{M}^{k})\odot\bm{d}_{IL}^{k}\right)$
 16
-  /* Meta-Update via BPTT (Outer-Level) */
+  /* Meta-Update via AutoDiff (Outer-Level) */
   $\ell_{\mathtt{meta}}(\bm{\omega}^{t})\leftarrow-\mathbb{E}_{f\sim p(f)}\left[\frac{\bar{f}(\mathbf{x}^{0})-\bar{f}(\mathbf{x}^{K}(\bm{\omega}^{t}))}{|\bar{f}(\mathbf{x}^{0})|+\epsilon}\right]$
-  // Cumulative Regret
+  // Normalized Improvement
 17
-18 $\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}\leftarrow\textnormal{{{AutoDiff}}}_{\text{BPTT}}(\ell_{\mathtt{meta}},\{\mathbf{x}^{k}\}_{k=0}^{K})$
-  $\bm{\omega}^{t+1}\leftarrow\bm{\omega}^{t}-\gamma\cdot\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}$
-  // Meta-Update
+ $\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}\leftarrow\textnormal{{{AutoDiff}}}(\ell_{\mathtt{meta}},\{\mathbf{x}^{k}\}_{k=0}^{K})$
+  // BPTT
+18  $\bm{\omega}^{t+1}\leftarrow\bm{\omega}^{t}-\gamma\cdot\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}$
 19
-20
-21Return $\mathbf{x}^{K}$, $\bm{\omega}^{T}$
+20Return $\mathbf{x}^{K}$, $\bm{\omega}^{T}$
 Algorithm 1 L2E: Differentiable Bilevel Evolution Strategy.
-The algorithm couples a neural fixed-point iteration with a dual-mode numerical operator: proxy gradients are injected as search directions during training, while inference remains purely evaluation-based.
+The algorithm couples a KM-style averaged neural update with dual-mode guidance: proxy gradients are injected during training, while inference remains evaluation-only.
 III-B Gradient-derived Evolutionary Composite Solver
-To instantiate the non-expansive operator $\mathcal{NU}(\cdot;\bm{\omega})$ introduced above, we propose a hybrid mechanism that fuses the global exploration capability of evolutionary algorithms with the local refinement power of gradient descent.
-This Gradient-Derived Evolutionary Composite (GEC) solver operates in a dual-mode fashion.
+To instantiate the proposed operator $\mathcal{NU}(\cdot;\bm{\omega})$ introduced above, we propose a hybrid mechanism that fuses the global exploration capability of evolutionary algorithms with the local refinement power of gradient descent.
+This GEC solver operates in a dual-mode fashion.
 III-B1 Learned evolutionary update
@@
 $\bm{d}_{IL}^{k}(\bm{\omega})=\mathcal{NU}(\bm{\mathbf{x}}^{k-1};\bm{\omega}),$ (4)
-where $\mathcal{NU}$ follows the KM-style fixed-point update defined in Eq. (2), composed of the learnable operator $\mathcal{O}_{\bm{\omega}}$.
+where $\mathcal{NU}$ follows the KM-style fixed-point update defined in Eq. (2), which is composed of the learnable operator $\mathcal{O}_{\bm{\omega}}$.
 III-B2 Proxy Gradient Update
 $\bm{d}_{OL}^{k}(\bm{\omega})=\bm{\mathbf{x}}^{k-1}-s_{k}\,\bm{P}_{\bm{\omega}}^{-1}\nabla_{\bm{\mathbf{x}}}\,\ell_{\mathtt{meta}}(\bm{\mathbf{x}}^{k-1},\bm{\omega}),$ (5)
-where $s_{k}$ denotes the inner step size and $\bm{P}_{\bm{\omega}}$ is a preconditioning matrix (e.g., identity, diagonal Hessian, or low-rank approximation).
+where $s_{k}$ denotes the inner step size, and $\bm{P}_{\bm{\omega}}$ represents a preconditioning matrix, such as an identity matrix, a diagonal Hessian, or a low-rank approximation.
 III-B3 Composite update with soft-gated fusion
-To seamlessly integrate these paths, we introduce an Adaptive Manifold Rectification mechanism.
 A dimension-wise gating mask $\mathbf{M}^{k}$ is computed based on the relative fitness improvement $\Delta f^{k}=f(\bm{d}_{OL}^{k})-f(\bm{d}_{IL}^{k})$, and apply a smooth gating function to produce a soft selection mask
 $\bm{M}^{k}=\sigma(-\Delta f^{k}/\tau),\quad\bm{M}^{k}\in(0,1)^{B\times N\times 1},$ (6)
-where $\sigma(\cdot)$ is the element-wise sigmoid function and $\tau$ is a temperature parameter controlling gating sharpness. The final update is a soft interpolation between the two candidates:
+where $\sigma(\cdot)$ is the element-wise sigmoid function and $\tau$ is a temperature parameter that controls the gating sharpness.
+The final update is computed as a soft interpolation between the two candidates:
 $\bm{\mathbf{x}}^{k}=\text{Proj}_{\mathcal{X},\bm{P}_{\bm{\omega}}}\left(\bm{M}^{k}\odot\bm{d}_{OL}^{k}+(1-\bm{M}^{k})\odot\bm{d}_{IL}^{k}\right),$ (7)
-where $\text{Proj}_{\mathcal{X},\bm{P}_{\bm{\omega}}}$ denotes a projection back to the feasible domain under the $\bm{P}_{\bm{\omega}}$-induced metric.
-This fusion mechanism balances exploitation and exploration in a differentiable, stability-preserving manner.
-III-B4 Meta-Objective Function
-To ensure robust meta-optimization across varying problem scales, we define $\ell_{\mathtt{meta}}$ based on the relative reduction of objective values.
-Let $\bar{f}(\mathbf{x})$ denote the mean objective value of a population $\mathbf{x}$.
-The meta-objective is formulated to minimize the negative expected normalized improvement over the task distribution $p(f)$:
+where $\text{Proj}_{\mathcal{X},\bm{P}_{\bm{\omega}}}$ denotes a projection back onto the feasible set under the $\bm{P}_{\bm{\omega}}$-induced metric.
+[^1] The subscript $(\mathcal{X},\mathbf{P}_{\bm{\omega}})$ indicates a metric projection onto the feasible set $\mathcal{X}$ under the $\mathbf{P}_{\bm{\omega}}$-induced norm: $\mathrm{Proj}_{\mathcal{X},\mathbf{P}_{\bm{\omega}}}(\mathbf{z}):=\arg\min_{\mathbf{x}\in\mathcal{X}}\ \|\mathbf{x}-\mathbf{z}\|_{\mathbf{P}_{\bm{\omega}}}^{2},\qquad\|\mathbf{v}\|_{\mathbf{P}_{\bm{\omega}}}^{2}:=\mathbf{v}^{\top}\mathbf{P}_{\bm{\omega}}\mathbf{v}.$ This fusion mechanism balances exploitation and exploration.
+It operates in a manner that is both differentiable and stability-preserving.
+III-B4 Meta-objective
+Let $\bar{f}(\mathbf{x})$ denote the mean objective value of a population $\mathbf{x}$. We define the meta-objective as the negative expected normalized improvement over the task distribution $p(f)$:
 $\ell_{\mathtt{meta}}(\bm{\omega})=-\mathbb{E}_{f\sim p(f)}\left[\frac{\bar{f}(\mathbf{x}^{0})-\bar{f}(\mathbf{x}^{K}(\bm{\omega}))}{|\bar{f}(\mathbf{x}^{0})|+\epsilon}\right],$ (8)
-where $\mathbf{x}^{K}(\bm{\omega})$ is the terminal population evolved on function $f$, and $\epsilon$ is a small constant for numerical stability.
-By minimizing this normalized loss, the framework is encouraged to maximize the proportion of cost reduction achieved within the unrolling horizon $K$, ensuring consistent optimization efficiency regardless of the absolute magnitude of the initial objective values.
-The complete execution protocol is formalized in Algorithm 1.
-This pseudocode unifies the theoretical framework into a coherent workflow: The inner loop (Lines 4-12) executes the composite update to enforce strict population contraction, while the outer loop (Lines 13-15) performs meta-optimization via BackPropagation Through Time (BPTT).
-By minimizing the derived meta-loss (Eq.
-8), the algorithm effectively aligns the operator’s fixed point with the target optimum, bridging the gap between theoretical design and practical implementation.
+where $\mathbf{x}^{K}(\bm{\omega})$ is the terminal population after $K$ unrolled updates on function $f$, and $\epsilon$ is a small constant for numerical stability.
+This normalization makes the meta-loss less sensitive to the absolute scale of objective values, enabling consistent training across tasks with different magnitudes.
+In Algorithm 1, we minimize $\ell_{\mathtt{meta}}$ via BPTT to update $\bm{\omega}$, thereby shaping the learned evolutionary dynamics to improve terminal performance under a fixed evaluation budget.
 III-C Structured Mamba Neural Operator
-To instantiate the evolution operator $\mathcal{O}_{evo}(\cdot;\omega)$, we introduce the Mamba-enhanced evolutionary operator, as illstrated in Fig.
-3.
-Unlike sequential SSMs, we implement this as a parallelized selective block to accommodate the permutation invariance of evolutionary populations, fusing local state refinement with global exploration.
+To instantiate the evolution operator $\mathcal{O}_{\mathrm{evo}}(\cdot;\bm{\omega})$, we adopt a structured Mamba-based operator implemented in a parallel selective form to respect the permutation-invariant nature of populations (Fig.
+3).
 III-C1 Parallel Mamba Block
-The core of this operator lies in the selection mechanism, where the SSM parameters are derived directly from the input to filter noise and propagate promising signals.
-We transform the raw population $\mathbf{x}\in\mathbb{R}^{B\times N\times D}$ and corresponding fitness values $\mathbf{f}$ into a compact embedding $\mathbf{E}$.
-Then we project $\mathbf{E}$ to generate the input-dependent discretization timescale $\bm{\Delta}$ and state transformation matrices $\mathbf{B},\mathbf{C}$.
-We decompose the operator into two parallel streams, a state expansion stream and an input gating stream, to approximate the discretized SSM dynamics without sequential scanning.
-The latent state $\mathbf{s}$ and the passthrough signal $\mathbf{u}$ (i.e., $\mathbf{u}=\phi_{in}(\mathbf{E})$) are computed as:
+The operator takes the population $\mathbf{x}\in\mathbb{R}^{B\times N\times D}$ and fitness values $\mathbf{f}$ as input and maps them to an embedding $\mathbf{E}$.
+We then generate input-dependent parameters, including the discretization timescale $\bm{\Delta}$ and state transformation matrices $\mathbf{B},\mathbf{C}$, and compute two parallel streams: a state-expanded representation and a passthrough signal.
+Specifically,
 $M_{\mathbf{s}}=(\bm{\Delta}\mathbf{B})\odot\mathbf{E},\,\,[\bm{\Delta},\mathbf{B},\mathbf{C}]=\text{Linear}_{proj}(\mathbf{E}).$ (9)
-Here, $M_{\mathbf{s}}$ represents the transformed evolutionary state modulated by the selective timescale $\bm{\Delta}$.
-The final output of the Mamba block is regulated by a learnable multiplicative gate $\mathbf{z}$, enabling dynamic information retention:
+The output is computed via a gated combination:
 $\small\mathbf{H}_{mamba}=\text{LN}\left(\mathbf{z}\odot M_{\mathbf{s}}+(1-\mathbf{z})\odot\mathbf{u}\right),\mathbf{z}=\sigma(\text{Linear}_{z}(\mathbf{E})).$ (10)
-This gated formulation ($\mathbf{z}\odot M_{\mathbf{s}}+(1-\mathbf{z})\odot\mathbf{u}$) mathematically mirrors the selective retention mechanism of Gated SSMs while ensuring stable gradient flow.
 III-C2 Hybrid Adaptive Fusion
-To complement the local refinement of Mamba with global population-level coordination, we integrate a parallel Multi-Head Self-Attention (MHSA) path: $\mathbf{H}_{attn}=\text{MHSA}(\mathbf{E})+\mathbf{E}.$ The final evolutionary proposal $\Delta_{evo}$ is synthesized via an adaptive routing mechanism that dynamically balances state exploitation and global exploration.
-The routing weights are determined by a lightweight non-linear router that processes population-level statistics.
-Specifically, a feature vector $\mathbf{s}_{pop}\in\mathbb{R}^{3}$ containing fitness standard deviation, fitness range, and population diversity is fed into an MLP to generate logits, which are then normalized:
+To incorporate population-level coordination, we add a parallel multi-head self-attention path $\mathbf{H}_{attn}=\text{MHSA}(\mathbf{E})+\mathbf{E}.$ The final evolutionary proposal is obtained by routing between the two paths using population statistics $\mathbf{s}_{pop}$:
 $\lambda_{ssm},\lambda_{attn}]=\text{Softmax}(\text{MLP}_{router}(\mathbf{s}_{pop})).$ (11)
-These adaptive coefficients then weigh the projected outputs of the two paths to form the final update proposal:
 $\small\Delta_{evo}=\lambda_{ssm}\cdot\text{Tanh}(\Psi_{m}(\mathbf{H}_{mamba}))+\lambda_{attn}\cdot\text{Tanh}(\Psi_{a}(\mathbf{H}_{attn})).$ (12)
-where $\Psi_{m}$ and $\Psi_{a}$ are projection heads, and the Tanh activation ensures the proposed updates are bounded within a controlled range for stability.
-Crucially, to strictly satisfy the theoretical Lipschitz continuity requirement of the GNU framework, we apply Spectral Normalization to all learnable linear transformations within both the Mamba and Attention blocks.
-By constraining the spectral norm of weight matrices, we ensure that the global Lipschitz constant of the operator $\mathcal{O}_{evo}$ is bounded (effectively $\leq 1$), thereby guaranteeing the contractive stability of the unrolled fixed-point iteration.
+Here, $\Psi_{m}$ and $\Psi_{a}$ denote projection heads.
+The Tanh activation bounds the update magnitude, serving as a stability-oriented design choice.
+Here we apply Spectral Normalization [48] to the learnable linear transformations in both paths to control update sensitivity.
+Together, spectral normalization and bounded updates act as stability-oriented regularization, making the learned updates less sensitive to input perturbations in practice.
 Figure 3: Architecture of the Structured Mamba Neural Operator.
 IV Theoretical Analysis
-In this section, we establish the theoretical foundation of L2E.
-Specifically, we demonstrate that L2E satisfies convergence theorems under a set of regularity assumptions.
-In contrast to heuristic meta-optimizers, which typically lack stability guarantees, the proposed non-expansive unrolling mechanism ensures a predictable convergence trajectory.
-The analysis proceeds in hierarchical steps.
-First, we quantify the finite-time convergence rates.
-Second, we establish the asymptotic consistency and bilevel optimality of the framework under stronger topological conditions.
-Due to space constraints, we provide detailed proofs for all propositions and theorems in the Supplementary Document.
+This section provides a stability-oriented theoretical analysis for L2E.
+We do not claim universal guarantees for arbitrary neural components.
+Instead, we clarify the sufficient conditions under which the unrolled operator iteration exhibits controlled dynamics and admits meaningful optimization bounds.
+Due to space constraints, detailed proofs are deferred to the Supplementary Document.
 IV-A Convergence Rate Analysis
@@
   1. (A1)
-    (Non-expansiveness) For any fixed $\bm{\omega}\in\Omega$, the learned composite operator $\mathcal{O}(\cdot;\bm{\omega})$ is non-expansive with respect to the population state $\mathbf{x}$: $\|\mathcal{O}(\mathbf{x};\bm{\omega})-\mathcal{O}(\mathbf{x}^{\prime};\bm{\omega})\|\leq\|\mathbf{x}-\mathbf{x}^{\prime}\|,\quad\forall\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}.$
+    (Operator regularity) For any fixed $\bm{\omega}\in\Omega$, the learned composite operator $\mathcal{O}(\cdot;\bm{\omega})$ is non-expansive with respect to the population state $\mathbf{x}$: $\|\mathcal{O}(\mathbf{x};\bm{\omega})-\mathcal{O}(\mathbf{x}^{\prime};\bm{\omega})\|\leq\|\mathbf{x}-\mathbf{x}^{\prime}\|,\ \forall\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}.$
   2. (A2)
-    (Smoothness) $\ell_{\mathtt{meta}}(\mathbf{x},\bm{\omega})$ is $L_{\mathbf{x}}$-smooth w.r.t. the population $\mathbf{x}$ and $L_{\bm{\omega}}$-smooth w.r.t. the parameters $\bm{\omega}$.
+    (Smoothness) $\ell_{\mathtt{meta}}(\mathbf{x},\bm{\omega})$ is $L_{\mathbf{x}}$-smooth with respect to the population $\mathbf{x}$ and $L_{\bm{\omega}}$-smooth with respect to the parameters $\bm{\omega}$.
   3. (A3)
@@
   4. (A4)
-    (Bounded domain) The initial population $\mathbf{x}^{0}$ and $\mathbf{x}^{*}(\bm{\omega})\in\text{Fix}(\mathcal{NU}(\cdot;\bm{\omega}))$ reside within a bounded region of diameter $D_{0}$: $\|\mathbf{x}^{0}-\mathbf{x}^{*}(\bm{\omega})\|\leq D_{0}$.
+    (Bounded domain) The initial population $\mathbf{x}^{0}$ and $\mathbf{x}^{*}(\bm{\omega})\in\text{Fix}(\mathcal{NU}(\cdot;\bm{\omega}))$ are contained within a bounded region of diameter $D_{0}$: $\|\mathbf{x}^{0}-\mathbf{x}^{*}(\bm{\omega})\|\leq D_{0}$.
 Based on these, we derive the approximation error of theinner loop and the convergence rate of the outer loop.
@@
 $\|\mathbf{x}^{K}(\bm{\omega})-\mathbf{x}^{*}(\bm{\omega})\|\leq(1-\alpha)^{K}D_{0},$ (13)
-where $\mathbf{x}^{*}(\bm{\omega})$ is the fixed point satisfying $\mathbf{x}^{*}=\mathcal{O}(\mathbf{x}^{*};\bm{\omega})$.
+where $\mathbf{x}^{*}(\bm{\omega})$ is the fixed point that satisfies $\mathbf{x}^{*}=\mathcal{O}(\mathbf{x}^{*};\bm{\omega})$.
 Theorem IV.1 (Convergence Rate).
 Under Assumptions (A1)-(A4), let the meta-parameters be updated via gradient descent: $\bm{\omega}^{t+1}=\bm{\omega}^{t}-\gamma\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}(\mathbf{x}^{K}(\bm{\omega}^{t});\bm{\omega}^{t})$.
-With step size $\gamma<\frac{1}{L_{\bm{\omega}}+L_{\mathbf{x}}C_{K}}$, the meta-gradient norm satisfies:
+Provided that the step size is $\gamma<\frac{1}{L_{\bm{\omega}}+L_{\mathbf{x}}C_{K}}$, the meta-gradient norm satisfies:
 $\min_{0\leq t<T}\|\nabla\ell_{\mathtt{meta}}(\bm{\omega}^{t})\|^{2}\leq\mathcal{O}\left(\frac{1}{T}+(1-\alpha)^{2K}\right).$ (14)
-Remark: The bound consists of an optimization error $\mathcal{O}(1/T)$ and an approximation bias $\mathcal{O}((1-\alpha)^{2K})$, confirming that L2E consistently improves the operator towards a stationary point of the meta-loss.
-IV-B Asymptotic Consistency and Bilevel Optimality
-To analyze the limiting behavior of the composite solver (Algorithm 1) and the exactness of the bilevel solution, we introduce stronger topological assumptions regarding the metric space induced by the preconditioner $\mathbf{P}_{\bm{\omega}}$.
+Remark: The bound decomposes into an optimization term $\mathcal{O}(1/T)$ and an approximation term $\mathcal{O}((1-\alpha)^{2K})$.
+This decomposition explicitly characterizes the influence of the unrolling depth $K$ and the averaging factor $\alpha$ on the meta-update quality under Assumptions (A1)-(A4).
+Moreover, the analysis aligns with established bilevel optimization literature, in which guarantees for convergence and consistency are derived from operator regularity and topological conditions [52, 36].
+IV-B Asymptotic Consistency under Fixed-point Constraints
+To interpret the limiting behavior of Algorithm 1 under the fixed-point constrained formulation, we consider stronger (idealized) topological assumptions.
+These assumptions facilitate an asymptotic consistency analysis.
+Specifically, this analysis connects the unrolled iteration to the fixed-point set of Eq.
+(1) and elucidates its relationship to the induced outer objective.
 Assumption IV.2.
@@
   2. (B2)
-    (Closedness) The graph $\mathrm{gph}(\mathcal{O}(\cdot;\bm{\omega})):=\{(\mathbf{u},\mathbf{v})\in\mathcal{X}\times\mathcal{X}\mid\mathbf{v}=\mathcal{O}(\mathbf{u};\bm{\omega})\}$, is closed.
+    (Closedness) The graph $\mathrm{gph}(\mathcal{O}(\cdot;\bm{\omega})):=\{(\mathbf{u},\mathbf{v})\in\mathcal{X}\times\mathcal{X}\mid\mathbf{v}=\mathcal{O}(\mathbf{u};\bm{\omega})\}$ is closed.
   3. (B3)
     (Compactness and Continuity) The sets $\Omega$ and $\mathcal{X}$ are compact.
-For any $\bm{\omega}$, $\text{Fix}(\mathcal{NU}(\cdot;\bm{\omega}))$ is non-empty.
-$\ell_{\mathtt{meta}}$ is continuous on $\mathcal{X}\times\Omega$, and for fixed $\bm{\omega}$, $\ell_{\mathtt{meta}}(\cdot;\bm{\omega})$ is $L_{\ell}$-smooth, convex, and lower-bounded.
-Under these structural conditions, we establish two fundamental properties: the asymptotic consistency of the inner solver and the optimality of the bilevel solution.
+Furthermore, for any $\bm{\omega}$, the set $\text{Fix}(\mathcal{NU}(\cdot;\bm{\omega}))$ is non-empty.
+Additionally, $\ell_{\mathtt{meta}}$ is continuous on $\mathcal{X}\times\Omega$.
+Finally, for a fixed $\bm{\omega}$, the function $\ell_{\mathtt{meta}}(\cdot;\bm{\omega})$ is $L_{\ell}$-smooth, convex, and lower-bounded.
 Theorem IV.2.
-Under Assumptions IV.2, we establish the following convergence properties for the L2E framework:
-(i) Asymptotic Consistency: Under Assumptions IV.2, let $\{\mathbf{x}^{k}(\bm{\omega})\}$ be the sequence generated by Algorithm 1.
-For any $\bm{\omega}\in\Omega$, provided the step size satisfies $s_{k}=\frac{\kappa}{k+1}$ with $0<\kappa<\frac{\lambda_{\min}(\mathbf{P}_{\bm{\omega}})}{L_{\mathbf{x}}}$, the sequence converges to the fixed-point set, i.e., $\lim_{k\to\infty}\mathrm{dist}(\mathbf{x}^{k}(\bm{\omega}),\mathrm{Fix}(\mathcal{NU}(\cdot;\bm{\omega})))=0$, and the meta-objective value converges to the optimum:
+Based on Assumption IV.2, we establish the following convergence properties for L2E:
+(i) Let $\{\mathbf{x}^{k}(\bm{\omega})\}$ denote the sequence generated by Algorithm 1.
+For any $\bm{\omega}\in\Omega$, assume the step size satisfies $s_{k}=\frac{\kappa}{k+1}$, where $0<\kappa<\frac{\lambda_{\min}(\mathbf{P}_{\bm{\omega}})}{L_{\mathbf{x}}}$.
+Under these conditions, the sequence converges to the fixed-point set, i.e., $\lim_{k\to\infty}\mathrm{dist}(\mathbf{x}^{k}(\bm{\omega}),\mathrm{Fix}(\mathcal{NU}(\cdot;\bm{\omega})))=0$.
+Consequently, the meta-objective value converges to the optimum:
 $\lim_{k\to\infty}\ell_{\mathtt{meta}}(\mathbf{x}^{k}(\bm{\omega});\bm{\omega})=\inf_{\mathbf{x}\in\mathrm{Fix}(\mathcal{NU}(\cdot;\bm{\omega}))}\ell_{\mathtt{meta}}(\mathbf{x};\bm{\omega}).$ (15)
-(ii) Bilevel Optimality: As the unrolling depth $K\to\infty$, let $\{(\mathbf{x}^{K}(\bm{\omega}^{K}),\bm{\omega}^{K})\}$ be the sequence of generated solutions.
-We prove that any limit point $(\bar{\mathbf{x}},\bar{\bm{\omega}})$ constitutes a valid solution to the fixed-point constrained bilevel problem (Eq.
-1), satisfying $\bar{\bm{\omega}}\in\arg\min_{\bm{\omega}\in\Omega}\varphi(\bm{\omega})$ and $\bar{\mathbf{x}}=\mathcal{NU}(\bar{\mathbf{x}};\bar{\bm{\omega}})$.
-Furthermore, the sequence of surrogate objectives converges to the true optimum, i.e., $\lim_{K\to\infty}\inf_{\bm{\omega}\in\Omega}\varphi_{K}(\bm{\omega})=\inf_{\bm{\omega}\in\Omega}\varphi(\bm{\omega})$.
+(ii) Let $\{(\mathbf{x}^{K}(\bm{\omega}^{K}),\bm{\omega}^{K})\}$ denote the sequence of generated solutions as the unrolling depth $K\to\infty$.
+We define the induced outer objective (value function) as $\varphi(\bm{\omega}):=\inf_{\mathbf{x}\in\mathrm{Fix}(\mathcal{NU}(\cdot;\bm{\omega}))}\ \ell_{\mathtt{meta}}(\mathbf{x};\bm{\omega}),$ $\varphi_{K}(\bm{\omega}):=\ell_{\mathtt{meta}}(\mathbf{x}^{K}(\bm{\omega});\bm{\omega}).$ We prove that any limit point $(\bar{\mathbf{x}},\bar{\bm{\omega}})$ is a valid solution to the fixed-point constrained bilevel problem (Eq.
+(1)).
+In particular, the limit point satisfies $\bar{\bm{\omega}}\in\arg\min_{\bm{\omega}\in\Omega}\varphi(\bm{\omega})$ and $\bar{\mathbf{x}}=\mathcal{NU}(\bar{\mathbf{x}};\bar{\bm{\omega}})$.
+Furthermore, the sequence of surrogate objectives converges to the optimum of the induced outer objective, i.e., $\lim_{K\to\infty}\inf_{\bm{\omega}\in\Omega}\varphi_{K}(\bm{\omega})=\inf_{\bm{\omega}\in\Omega}\varphi(\bm{\omega})$.
 V Experimental Study
@@
 We evaluate the optimization efficiency and zero-shot generalization capability of L2E using a hierarchical benchmark suite.
-This suite ranges from synthetic functions to real-world robot control tasks.
-We structure the evaluation protocol into three categories based on the relationship between the training and testing distributions:
-  • Synthetic Benchmarks (In-Distribution): We utilize BBOB-10D, BBOB-30D [54], and LSGO-1000D [55] as the primary testbeds.
-These datasets constitute both the training sets and the in-distribution test sets.
-Specifically, the BBOB tasks evaluate basic optimization capabilities with budgets of $2\times 10^{4}$ (10D) and $5\times 10^{4}$ (30D).
-In contrast, LSGO assesses scalability in high-dimensional regimes ($D\approx 1000$) with a budget of $3\times 10^{6}$.
-  • Out-of-Distribution (OOD) Generalization: We select two benchmarks to assess the transferability of L2E across disparate domains.
-First, we employ the UAV benchmark [56], which comprises 56 real-world path planning instances.
-Here, we test a model trained exclusively on BBOB-30D under a constrained budget of $2.5\times 10^{2}$.
-Second, we utilize BBOB-surrogate-10D, i.e., a synthetic variant derived from surrogate model landscapes.
-This benchmark tests robustness against landscape shifts and approximation errors.
-In this case, we apply the model trained on standard BBOB-10D directly without fine-tuning.
-  • Neuroevolution for Robot Control (Flexibility Verification): We extend the evaluation to the NE benchmark [57], which consists of high-dimensional robot control policy search tasks ($D>1000$).
-We use this suite specifically to benchmark L2E against reinforcement learning methods.
-This comparison assesses the flexibility of the framework in control scenarios, given a maximum evaluation budget of $2,500$.
+This suite encompasses tasks ranging from synthetic functions to real-world robot control.
+Regarding hyperparameters, we set the relaxation factor $\alpha=0.9$, the gating temperature $\tau=1$, the unrolling horizon $K=200$, and the population size $B=100$, unless otherwise specified.
+For spectral normalization, we use the default coefficient $1.0$ for all linear layers.
+We conduct all experiments on a single NVIDIA RTX A6000 GPU (PyTorch 2.0.1).
+Each task is evaluated over 10 independent random seeds, and we report both the mean and standard deviation.
 TABLE I: Numerical performance evaluation (mean $\pm$ std) of representative algorithms on the BBOB-10D benchmark.
@@
 ($\pm$2.988E+01) ($\pm$6.491E+00) ($\pm$7.657E+00) ($\pm$1.685E+02) ($\pm$4.851E+02) ($\pm$2.196E+04) ($\pm$6.461E+00) ($\pm$5.398E+05)
-CMAES 1.514E+03 1.838E+04 8.980E+01 1.831E+04 2.004E+02 1.612E+06 6.288E+01 7.855E+07
-($\pm$4.183E+00) ($\pm$1.386E+02) ($\pm$3.532E-01) ($\pm$3.221E+02) ($\pm$6.800E+00) ($\pm$6.243E+03) ($\pm$1.963E-01) ($\pm$3.370E+05)
 SAHLPSO 1.424E+02 1.167E+02 2.402E+01 9.622E+02 1.070E+03 4.344E+04 1.141E+02 7.618E+06
@@
 […]
-Method Sharp_Rid Different_Pow Schaffers_hig Composite_Gri Schwefel Gallagher_21P Katsuura Lunacek_bi_Ras
-PSO 9.431E+02 7.173E+00 2.411E+01 4.266E+00 3.235E+02 6.513E+01 1.209E+00 1.166E+02
-($\pm$1.627E+02) ($\pm$3.197E+00) ($\pm$9.562E-01) ($\pm$6.784E-01) ($\pm$3.334E+02) ($\pm$5.339E+00) ($\pm$2.412E-01) ($\pm$7.371E+00)
-DE 2.583E+02 2.088E+00 1.081E+01 4.509E+00 2.413E+00 1.420E+01 2.733E+00 8.125E+01
-($\pm$2.861E+01) ($\pm$1.005E+00) ($\pm$8.182E-01) ($\pm$2.996E-01) ($\pm$9.219E-02) ($\pm$1.058E+01) ($\pm$7.799E-01) ($\pm$3.929E+00)
-CMAES 1.309E+03 1.309E+01 4.228E+01 3.101E+00 1.671E+03 7.707E+01 5.306E-01 6.387E+01
-($\pm$1.568E+00) ($\pm$4.955E-02) ($\pm$5.127E-03) ($\pm$5.885E-01) ($\pm$4.225E+01) ($\pm$1.659E-02) ($\pm$6.150E-02) ($\pm$4.391E-02)
-SAHLPSO 3.679E+02 1.316E+00 1.229E+01 5.759E+00 3.942E+01 1.675E+01 1.577E+00 9.341E+01
-($\pm$2.165E+01) ($\pm$1.267E+00) ($\pm$5.473E+00) ($\pm$5.528E-01) ($\pm$3.132E+01) ($\pm$5.020E+00) ($\pm$5.575E-01) ($\pm$2.748E+00)
-RNNOPT 1.798E+03 1.069E+01 5.223E+01 2.056E+00 8.060E+03 8.486E+01 3.440E+00 1.221E+02
-($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00)
-DEDQN 1.057E+03 1.047E+01 3.073E+01 1.127E+01 4.618E+03 5.834E+01 3.613E+00 1.480E+02
-($\pm$1.827E+02) ($\pm$3.672E+00) ($\pm$4.258E+00) ($\pm$1.689E+00) ($\pm$8.773E+02) ($\pm$9.502E+00) ($\pm$5.164E-01) ($\pm$2.998E+01)
-LES 3.274E+03 1.505E+03 7.211E+02 1.700E+03 2.104E+03 2.454E+03 6.017E+02 1.270E+03
-($\pm$6.658E+01) ($\pm$3.239E-01) ($\pm$6.131E-01) ($\pm$3.196E-02) ($\pm$2.379E-01) ($\pm$1.005E+01) ($\pm$1.436E-01) ($\pm$2.917E+00)
-LGA 9.683E+02 1.663E+01 2.613E+01 1.322E+01 1.234E+04 4.875E+01 1.618E+00 1.778E+02
-($\pm$2.034E+02) ($\pm$5.850E+00) ($\pm$4.232E+00) ($\pm$5.755E+00) ($\pm$3.810E+03) ($\pm$1.666E+01) ($\pm$0.000E+00) ($\pm$1.977E+01)
-GLHF 1.111E+03 1.530E+01 2.792E+01 9.223E+00 4.686E+03 4.876E+01 3.299E+00 1.499E+02
-($\pm$2.052E+02) ($\pm$8.705E-01) ($\pm$3.296E+00) ($\pm$1.221E+00) ($\pm$1.331E+03) ($\pm$1.684E+01) ($\pm$4.095E-01) ($\pm$2.572E+01)
-B2OPT 4.721E+02 4.665E+00 1.176E+01 1.633E+00 2.877E+00 2.987E+01 2.293E+00 7.342E+01
-($\pm$6.671E+01) ($\pm$1.035E+00) ($\pm$2.675E+00) ($\pm$1.135E-01) ($\pm$3.331E-01) ($\pm$1.920E+01) ($\pm$4.472E-02) ($\pm$5.952E+00)
-[…]
-[…]
-[…]
-[…]
 indicates ours variant enhanced with the proposed Mamba-based neural operator.
+Due to space limitations, only a subset is shown here; the full table is provided in the Supplementary Material (Table S.I).
+  • Synthetic Benchmarks (In-Distribution): We utilize BBOB-10D, BBOB-30D [20], and LSGO-1000D [33] as the primary testbeds.
+These benchmarks represent the same function families utilized during meta-training, although we conduct the evaluation on held-out runs or instances under identical protocols.
+Specifically, BBOB assesses general optimization capability with budgets of $2\times 10^{4}$ (10D) and $5\times 10^{4}$ (30D), whereas LSGO evaluates scalability in high-dimensional regimes ($D\approx 1000$) with a budget of $3\times 10^{6}$.
+  • Out-of-Distribution (OOD) Evaluation: We incorporate two benchmarks to assess robustness beyond the training distribution.
+First, we employ the UAV benchmark [53], which consists of 56 real-world path planning instances.
+Specifically, we directly evaluate a model trained on BBOB-30D under a constrained budget of $2.5\times 10^{2}$ without fine-tuning.
+Second, we use BBOB-surrogate-10D, a synthetic variant derived from surrogate-induced landscapes, to test robustness to landscape shifts and approximation mismatch; we apply the model trained on BBOB-10D directly.
+  • Neuroevolution for Robot Control: We additionally evaluate performance on the NE benchmark [22], which comprises high-dimensional policy search tasks ($D>1000$) for robot control.
+This suite assesses the applicability of L2E in control scenarios and benchmarks it against representative reinforcement learning methods under a maximum evaluation budget of 2,5002{,}500.
 TABLE II: Numerical performance evaluation (mean $\pm$ std) of representative algorithms on the BBOB-30D benchmark..
@@
 ($\pm$1.715E+01) ($\pm$7.413E+02) ($\pm$3.713E+01) ($\pm$7.355E+02) ($\pm$5.186E+02) ($\pm$1.613E+05) ($\pm$1.608E+01) ($\pm$3.474E+06)
-CMAES 4.236E+03 1.409E+05 1.435E+03 4.085E+04 1.850E+02 2.423E+06 1.732E+02 1.689E+08
-($\pm$4.880E+01) ($\pm$1.602E+03) ($\pm$1.935E+01) ($\pm$1.404E+02) ($\pm$3.947E+00) ($\pm$1.120E+04) ($\pm$6.706E-01) ($\pm$1.487E+06)
 SAHLPSO 9.574E+02 5.158E+04 4.115E+02 2.687E+04 9.397E+03 1.296E+06 2.760E+02 1.124E+08
@@
 […]
-Method Sharp_Rid Different_Pow Schaffers_hig Composite_Gri Schwefel Gallagher_21P Katsuura Lunacek_bi_Ras
-PSO 1.360E+03 1.928E+01 2.751E+01 5.665E+00 2.016E+01 6.659E+01 3.208E+00 3.461E+02
-($\pm$1.372E+02) ($\pm$3.804E+00) ($\pm$4.309E+00) ($\pm$2.647E-01) ($\pm$1.457E+01) ($\pm$2.572E+00) ($\pm$2.201E-01) ($\pm$2.614E+01)
-DE 7.012E+02 8.848E+00 1.536E+01 6.595E+00 2.439E+02 1.083E+01 3.000E+00 3.043E+02
-($\pm$3.727E+01) ($\pm$7.387E-01) ($\pm$5.156E+00) ($\pm$1.962E-01) ($\pm$1.181E+02) ($\pm$7.941E-01) ($\pm$1.823E-01) ($\pm$3.679E+01)
-CMAES 1.485E+03 2.591E+01 4.616E+01 4.878E+00 3.021E+00 7.824E+01 4.913E-01 2.853E+02
-($\pm$1.474E+00) ($\pm$9.453E-02) ($\pm$4.374E-02) ($\pm$1.767E-01) ($\pm$6.518E-02) ($\pm$7.518E-02) ($\pm$1.826E-01) ($\pm$1.699E+00)
-SAHLPSO 1.500E+03 1.707E+01 3.268E+01 8.649E+00 4.673E+03 4.442E+01 3.192E+00 4.744E+02
-($\pm$3.082E+01) ($\pm$8.338E+00) ($\pm$4.166E+00) ($\pm$1.578E+00) ($\pm$1.368E+03) ($\pm$8.733E+00) ($\pm$7.748E-01) ($\pm$2.208E+01)
-RNNOPT 2.314E+03 7.363E+01 7.426E+01 6.430E+00 2.606E+04 8.447E+01 7.372E+00 4.541E+02
-($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00) ($\pm$0.000E+00)
-DEDQN 2.360E+03 4.823E+01 4.896E+01 2.184E+01 6.849E+04 7.642E+01 4.808E+00 8.033E+02
-($\pm$1.659E+02) ($\pm$2.662E+00) ($\pm$5.639E+00) ($\pm$1.482E+00) ($\pm$1.161E+04) ($\pm$3.011E+00) ($\pm$3.796E-01) ($\pm$9.052E+01)
-LES 2.895E+03 1.310E+02 1.341E+03 1.300E+03 7.007E+03 1.180E+03 1.303E+03 1.749E+03
-($\pm$3.334E+01) ($\pm$1.310E+00) ($\pm$2.615E+00) ($\pm$4.597E-04) ($\pm$1.338E+03) ($\pm$1.277E+00) ($\pm$2.324E-01) ($\pm$1.423E+01)
-LGA 2.821E+03 9.115E+01 6.749E+01 2.159E+01 9.618E+04 8.395E+01 3.231E+00 9.121E+02
-($\pm$1.272E+02) ($\pm$2.051E+01) ($\pm$5.473E+00) ($\pm$3.009E+00) ($\pm$1.232E+04) ($\pm$1.719E+00) ($\pm$2.440E-01) ($\pm$1.066E+02)
-GLHF 2.611E+03 6.510E+01 5.337E+01 2.541E+01 7.450E+04 8.305E+01 6.515E+00 8.946E+02
-($\pm$1.152E+02) ($\pm$5.242E+00) ($\pm$5.011E+00) ($\pm$1.609E+00) ($\pm$4.114E+03) ($\pm$1.280E+00) ($\pm$1.767E-01) ($\pm$3.887E+01)
-B2OPT 1.393E+03 1.921E+01 3.485E+01 2.156E+00 1.671E+03 7.715E+01 4.219E+00 3.267E+02
-($\pm$1.978E+02) ($\pm$3.729E-01) ($\pm$1.291E+00) ($\pm$2.386E-01) ($\pm$8.537E+02) ($\pm$3.433E+00) ($\pm$2.024E-01) ($\pm$5.097E+00)
-[…]
-[…]
-[…]
-[…]
 indicates ours variant enhanced with the proposed Mamba-based neural operator.
+Due to space limitations, only a subset is shown here; the full table is provided in the Supplementary Material (Table S.II).
 Compared Methods.
-We conduct a comparative analysis against representative algorithms to evaluate the effectiveness of the proposed L2E framework.
-This set encompasses both classical heuristics and learning-based optimizers.
-To situate our non-expansive unrolling paradigm within the broader L2O landscape, we incorporate two categories of learned algorithms.
-The first category consists of meta-learning-based neuro-evolution strategies, i.e., DEDQN [58], LES [28], and LGA [59].
-The second category includes control-based solution generation models, i.e., RNNOPT [60], GLHF [61], and B2OPT [23].
-Additionally, we employ established traditional algorithms, i.e., CMA-ES [11], DE [62], PSO [31], and SAHLPSO [63], which serve as performance anchors.
-We re-evaluate all competing methods within a unified experimental protocol and identical computational budgets, thereby ensuring a consistent assessment.
-Finally, we conduct all experiments on a single NVIDIA RTX A6000 GPU (PyTorch 2.0.1) and evaluate each task over 10 independent random seeds to report the mean and standard deviation.
-V-B Performance on Numerical Benchmarks
-We first examine the optimization capability of L2E on numerical landscapes where the meta-training and testing distributions align.
-This assessment covers two critical dimensions: precision on standard numerical problems (BBOB) and scalability to high dimensions (LSGO).
-V-B1 Precision and Stability on Standard Benchmarks
-Table I and Table II report the comparative results on BBOB-10D and BBOB-30D, respectively. L2E demonstrates a substantial advantage over state-of-the-art methods:
-  • Performance Superiority and Architectural Efficacy: L2E establishes a consistent performance advantage.
-Specifically, it achieves the lowest mean error on 31 out of 32 test instances.
-On ill-conditioned and multimodal functions (e.g., Rosenbrock_rot, Lunacek_bi_Ras), L2E substantially outperforms strong baselines, such as GLHF and B2OPT.
-We attribute this escape capability to the non-expansive unrolling mechanism, which enforces a strict contractive constraint to prevent divergence in high-curvature valleys.
-Furthermore, the comparison between variants reveals that the Mamba-enhanced model (Ours) consistently improves upon the base model (Ours), particularly on non-separable landscapes (e.g., Ellipsoidal_hc).
-This result confirms that the structured state-space model offers a superior inductive bias for capturing complex variable couplings, in contrast to the generic attention or coordinate-wise mechanisms used in the standard version.
-  • Convergence Trajectory Analysis: Figure 4 and Figure 5 illustrate the convergence behavior.
-L2E (red curves) exhibits a steeper initial descent and tighter confidence intervals.
-On unimodal functions (Bent Cigar), the model rapidly saturates to the optimum.
-This behavior demonstrates the effective exploitation of gradient-like information.
-Conversely, on multimodal tasks (Gallagher 21Peaks), L2E maintains steady progress without early stagnation.
-This balance confirms the efficacy of the Elite-Mask Fusion, which adaptively arbitrates between global exploration and local gradient refinement.
-V-B2 Scalability to High-Dimensional Regimes
-The results on the 1000D LSGO benchmark (Table III) further highlight the structural advantages of L2E.
-Specifically, L2E ranks first across all six representative high-dimensional functions.
-On the complex Shifted_Rastrigin, the model achieves a lower objective value, a result that surpasses LGA and GLHF.
-Existing Transformer-based optimizers (e.g., GLHF) encounter substantial bottlenecks in 1000D regimes due to the quadratic complexity of the self-attention mechanism.
-In contrast, our Mamba-based operator maintains linear complexity, which facilitates stable gradient propagation during meta-training.
-The ECDF curves in Fig.
-6 and boxplots in Fig.
-7 illustrate this distinction.
-While RNNOPT and DEDQN exhibit flatlining trajectories, which indicate gradient vanishing or saturation, L2E maintains a continuous descent.
-This behavior demonstrates the capability of L2E to navigate high-dimensional manifolds efficiently.
+We conduct a comparative analysis against representative algorithms to evaluate the effectiveness of the proposed L2E framework, covering both classical heuristics and learning-based optimizers.
+To situate our stability-inducing unrolling framework within the broader L2O landscape, we include two categories of learned optimizers.
+The first category consists of meta-learning-based neuro-evolution strategies, including DEDQN [56], LES [27], and LGA [26].
+The second category includes control-based solution generation models, including RNNOPT [59], GLHF [31], and B2OPT [32].
+In addition, we include representative classical population-based baselines, including DE [54], PSO [25], and SAHLPSO [58], to provide reference performance under the same evaluation budgets.
+We re-evaluate all competing methods under a unified experimental protocol and identical computational budgets.
 Attractive Sector
@@
 Figure 4: Log-scaled convergence trajectories of our method and representative methods on the first 8 benchmark tasks from the BBOB-10D suite [20].
-The x-axis denotes the number of function evaluations (FEs), while the y-axis reports the logarithm of the objective value (log cost), where lower values indicate superior performance.
-For clarity, only 8 representative tasks are visualized; Complete results on all 16 functions are provided in the supplementary material (i.e., Fig S1).
-V-C Generalization to Real-World UAV Planning
-A core contribution of L2E is its capability to generalize to unseen, real-world tasks without fine-tuning.
-We validate this Sim-to-Real transfer capability using the UAV Path Planning benchmark.
-Although the model is trained exclusively on synthetic BBOB-30D functions, L2E generalizes robustly to the 56 distinct UAV planning tasks (30D).
-As depicted in the radar plots of Fig.
-8, L2E achieves the highest normalized performance across all seven terrain groups.
-The improvement margins over the second-best baseline range from 3.4% to 14.2%.
-This performance advantage stems from the proposed transfer mechanism.
-In contrast to reinforcement learning-based methods (e.g., DEDQN), which risk overfitting to specific landscape features encountered during training, L2E learns a generic descent operator that is governed by fixed-point properties.
-Consequently, the optimizer perceives the irregular constraints and noise profiles of the UAV simulator as intrinsic properties of the black-box function.
-This perception allows it to apply the learned contractive rules to navigate feasible regions.
-V-D Neuroevolution for Robot Control
-Finally, we assess the applicability of the framework to high-dimensional policy search tasks in the Neuroevolution (NE) benchmark [57].
-This assessment represents a substantial domain shift from function optimization to control policy learning.
-The benchmark consists of six realistic robotic tasks across two domains, i.e., ant and humanoid.
-Specifically, each domain contains three complexity variants (ant-3, ant-4, ant-6; and humanoid-3, humanoid-4, humanoid-5).
-We benchmark the proposed method against GLEET [64], i.e., a specialized baseline that formulates neuroevolution within a reinforcement learning framework.
-GLEET trains an actor-critic agent to evolve population-level solutions via environmental interactions.
-We conduct this comparison to demonstrate that the learned optimizer can replace heuristic operators to optimize actor-critic evolutionary control policies.
-Fig.
-9 presents the comparative results, utilizing a normalized metric to enable intuitive visual comparison across tasks with distinct reward scales.
-The vertical axis represents deviations from a task-specific mean, rescaled by a common exponential factor: $y_{\text{plot}}=\frac{y-\text{mean}}{10^{\text{exp}}}.$ This transformation highlights the convergence dynamics around a central baseline, and it preserves cross-task comparability.
-On representative mid-difficulty tasks, such as ant-4 and humanoid-4, L2E exhibits steeper convergence trajectories and lower terminal costs than the baseline.
-Furthermore, the aggregated boxplots (Right) reveal that L2E achieves lower final objective values with consistently reduced variance across all six tasks.
-This result suggests that replacing heuristic population updates with the learned non-expansive operator stabilizes the policy learning process, thereby mitigating the instability typically associated with high-variance reward signals in RL environments.
+Complete results on all 16 functions are provided in the supplementary material (i.e., Fig S1).
+V-B Performance on Numerical Benchmarks
+We first examine the optimization capability of L2E on numerical landscapes where the meta-training and testing distributions align.
+This assessment focuses on two aspects: performance on standard numerical benchmarks (BBOB) and scalability to high dimensions (LSGO).
+V-B1 Precision and Stability on Standard Benchmarks
+Table I and Table II report the comparative results on BBOB-10D and BBOB-30D, respectively. Overall, L2E achieves consistently strong performance across a wide range of test instances:
+  • Performance and Architectural Efficacy: L2E attains the lowest mean error on 31 out of 32 test instances.
+Specifically, L2E outperforms established learned baselines, such as GLHF and B2OPT, on ill-conditioned and multimodal functions (e.g., Rosenbrock_rot, Lunacek_bi_Ras).
+We attribute these gains to the stability-inducing unrolling design and the composite update structure, which promote controlled optimization dynamics in high-curvature regions.
+Furthermore, the Mamba-based variant (Ours) consistently improves upon the base model (Ours), particularly on non-separable landscapes (e.g., Ellipsoidal_hc).
+This finding suggests that the structured state-space operator provides an effective inductive bias for capturing population-wise variable couplings.
+  • Trajectory and Stability Analysis: L2E demonstrates robust convergence behaviors, as visualized in Figure 4 and Figure 5.
+Specifically, the method exhibits faster early-stage progress with tighter confidence intervals.
+On unimodal functions (e.g., Bent Cigar), L2E quickly reaches a low-error regime, which indicates effective exploitation when the landscape is smooth.
+In contrast, on multimodal tasks (e.g., Gallagher 21Peaks), L2E maintains steady improvement without early stagnation.
+This sustained progress reflects the benefit of the Elite-Mask Fusion, which adaptively balances learned evolutionary proposals and proxy numerical guidance.
+V-B2 Scalability to High-Dimensional Regimes
+Results on the 1000D LSGO benchmark (Table III) further demonstrate the scalability of L2E.
+L2E ranks first across all six representative high-dimensional functions.
+Figure 6 (ECDF) and Fig.
+7 (boxplots) illustrate the scalability of L2E in high-dimensional regimes.
+In contrast to existing learned optimizers, L2E exhibits consistent progress and reduced performance variance in the 1000D regime.
+This robustness stems from the structured operator design and the regularization strategies, i.e., bounded updates and soft-gated fusion.
+Consequently, these mechanisms promote controlled dynamics during optimization in high-dimensional spaces.
 Attractive Sector
@@
 Figure 5: Log-scaled convergence trajectories of our method and representative methods on the first 8 benchmark tasks from the BBOB-30D suite [20].
-The x-axis denotes the number of function evaluations (FEs), while the y-axis reports the logarithm of the objective value (log cost), where lower values indicate superior performance.
-For clarity, only 8 representative tasks are visualized; Complete results on all 16 functions are provided in the supplementary material (i.e., Fig S2).
+Complete results on all 16 functions are provided in the supplementary material (i.e., Fig S2).
 TABLE III: High-Dimensional numerical evaluation (mean $\pm$ std) of representative algorithms on the LSGO-1000D.
@@
 […]
-VI Ablation Analysis and Mechanistic Insights
-To disentangle the contribution of each algorithmic component and justify our design choices, we conduct a systematic ablation study on the BBOB-10D suite.
-This analysis is structured into two parts: (1) an investigation of the neural parameterization strategy for the evolution operator, specifically Shared versus Unshared Weights; and (2) a dissection of the combined efficacy of the core modules, i.e., ProxyGrad, SoftGate, and Mamba.
-VI-A Evolutionary Curriculum via Unshared Parameterization
-A pivotal architectural decision in deep unrolling is the strategy for sharing the meta-parameters $\mathcal{O}_{\mathrm{evo}}(\cdot;\bm{\omega})$ of the evolution operator $\mathcal{O}_{\mathrm{evo}}$ across time steps $k=1\dots K$.
-  • Shared ($\bm{\omega}_{\text{shared}}$): This strategy ties a single set of weights across all unrolled blocks, which forces the operator to learn a time-invariant update rule.
-  • Unshared ($\bm{\omega}^{k}$): This strategy allocates independent weights for each block, which allows the operator to evolve distinct behaviors at different stages of the optimization process.
 Rot. Ackley
@@
 Complete results on all 6 functions are provided in the supplementary material (i.e., Fig S8).
+V-C Generalization to Real-World UAV Planning
+We evaluate the robustness of L2E under distribution shifts using the UAV Path Planning benchmark.
+Although L2E is meta-trained exclusively on synthetic BBOB-30D functions, it is directly applied to 56 UAV planning instances (30D) without fine-tuning.
+As shown by the radar plots in Fig.
+8, L2E achieves the best normalized performance across all seven terrain groups, with relative gains over the second-best baseline ranging from 3.4% to 14.2%.
+We attribute this empirical transferability to the stability-inducing unrolling formulation and the structured composite updates in L2E.
+Compared with policy-style learned optimizers that may become brittle when the constraint patterns and noise characteristics differ from training, L2E relies on structured population updates with stability-oriented regularization (e.g., bounded proposals and soft-gated numerical guidance).
+These design choices promote controlled dynamics under the strict evaluation budget, enabling more reliable progress on irregular UAV landscapes.
+V-D Neuroevolution for Robot Control
 Rot. Ackley
@@
 The percentage value marked in each subplot denotes our method’s improvement margin over the second-best baseline.
+Finally, we evaluate the applicability of L2E to high-dimensional policy search in the Neuroevolution (NE) benchmark [22].
+This setting introduces a distinct distribution shift from synthetic function optimization to control-oriented policy search.
+The benchmark includes six robotic tasks across two domains, ant and humanoid, where each domain comprises three complexity variants (ant-3/4/6 and humanoid-3/4/5).
+We compare L2E against GLEET [43], a specialized baseline that frames neuroevolution within an actor–critic reinforcement learning pipeline.
+This comparison tests whether a learned evolutionary optimizer can serve as a drop-in alternative to handcrafted population update heuristics in control scenarios.
+Figure 9 reports normalized learning curves to enable a fair visual comparison across tasks with different reward scales.
+The normalization is defined as $y_{\text{plot}}=\frac{y-\text{mean}}{10^{\text{exp}}}.$ On representative mid-difficulty tasks (e.g., ant-4 and humanoid-4), L2E exhibits faster progress and improved final performance compared with the baseline.
+Moreover, the aggregated boxplots (Right) indicate that L2E achieves better terminal objective values with consistently reduced variance across all six tasks.
+These results suggest that the stability-inducing unrolling formulation and structured composite updates in L2E promote controlled policy search dynamics under noisy and high-variance feedback.
+Consequently, L2E improves robustness in neuroevolution-based control.
+VI Ablation Analysis and Mechanistic Insights
+To quantify the contribution of each component and motivate the design choices, we conduct systematic ablations on BBOB-10D.
+We focus on two questions: (1) how different parameterization strategies for the unrolled evolutionary operator affect optimization dynamics (shared vs.
+unshared weights across steps), and (2) how the core modules(i.e., ProxyGrad, SoftGate, and Mamba) influence performance and stability.
+VI-A Shared vs. Unshared Unrolling Parameters
+A pivotal design choice in neural unrolling determines whether to share the meta-parameters of the evolution operator $\mathcal{O}_{\mathrm{evo}}(\cdot;\bm{\omega})$ across unrolling steps $k=1,\dots,K$.
+This distinction establishes two primary configurations:
+  • Shared ($\bm{\omega}_{\text{shared}}$): A single set of weights is utilized across all unrolled blocks, which enforces a time-invariant update rule.
+  • Unshared ($\bm{\omega}^{k}$): Independent weights are assigned to distinct blocks, thereby enabling step-dependent behaviors that capture different phases of the optimization trajectory.
 Figure 9: Comparison results on the Neuroevolution for Robot Control benchmark [22], comprising six high-dimensional robot control tasks from two domains: ant (ant-3, ant-4, ant-6) and humanoid (humanoid-3, humanoid-4, humanoid-5).
@@
 The right plot shows boxplots of final objective values (log scale) aggregated across all six tasks.
-VI-A1 Statistical Superiority of Unshared Design
-The unshared parameterization strategy consistently outperforms the Shared baseline across all metrics.
-Table IV presents the aggregate performance across varying evaluation budgets.
-This advantage amplifies as the trajectory length increases.
-Specifically, at $\mathrm{MaxFEs}=20{,}000$ ($\mathrm{EMS}=200$), the Unshared model reduces the mean error by approximately 96% compared to the Shared version.
-This empirical gap suggests that a static operator is insufficient to handle the dynamic requirements of a complete optimization lifecycle.
-Conversely, the allocation of unique $\bm{\omega}$ to each block enables stage-wise specialization of the update dynamics.
-This capability is increasingly beneficial as $\mathrm{EMS}$ grows and the learning trajectory lengthens.
-Consequently, our final model adopts the separate per-block meta-parameters design for $\mathcal{O}_{\mathrm{evo}}$.
-Although this incurs a marginally higher training cost, it yields superior accuracy and stability under practical evaluation budgets.
 TABLE IV: Ablation on Evolution Neural Block weight sharing on BBOB-10D.
@@
 S9 and S10).
+VI-A1 Statistical Advantage of the Unshared Design
+Across evaluation budgets, the unshared parameterization achieves consistently better performance than the shared-weight baseline.
+Table IV summarizes the aggregated results under different evaluation budgets, where the performance gap becomes more pronounced as the unrolling horizon increases.
+In particular, at $\mathrm{MaxFEs}=20{,}000$ ($\mathrm{EMS}=200$), the unshared model reduces the mean error by approximately 96% relative to the shared variant.
+These results suggest that a time-invariant update rule can be insufficient to accommodate the evolving needs of the optimization trajectory.
+By assigning step-specific meta-parameters $\bm{\omega}^{k}$, the unshared design enables stage-dependent update behaviors (e.g., exploration-oriented early steps and refinement-oriented later steps), which becomes increasingly beneficial as $\mathrm{EMS}$ grows.
+Therefore, we adopt separate per-block meta-parameters for $\mathcal{O}_{\mathrm{evo}}$ in the final model.
+Although this choice increases training cost moderately, it yields better accuracy and more stable behavior under practical evaluation budgets.
 VI-A2 Stage-wise Specialization
-Visual analysis confirms that the unshared parameterization enables an implicit evolutionary curriculum.
-Fig.
-10 provides a visual explanation of this phenomenon.
-In the early stage, i.e., exploration, both configurations descend rapidly.
-However, in the late stage, i.e., exploitation, the Shared model (brown curves) flattens out prematurely.
-It functions as a stationary operator, which prevents mode switching.
-In contrast, the unshared model (red/purple curves) continues to refine the solution.
-The unshared parameterization effectively learns an implicit evolutionary curriculum.
-Specifically, early blocks specialize in global exploration to escape local basins, whereas later blocks shift towards precise local refinement.
-This time-dependent specialization is pivotal for high-precision convergence in black-box optimization.
+Figure 10 provides qualitative evidence that unshared parameterization enables stage-dependent update behaviors along the unrolled trajectory.
+In the early phase (exploration), both the shared and unshared designs make rapid progress.
+However, in the late phase (refinement), the shared-weight variant (brown curves) tends to plateau earlier, indicating limited flexibility to adapt its update behavior as the search proceeds.
+In contrast, the unshared variant (red/purple curves) continues to refine the solutions and achieves better late-stage improvements.
+Overall, this visualization suggests that step-specific meta-parameters $\bm{\omega}^{k}$ allow different unrolling blocks to specialize in different roles with earlier blocks favoring broader exploration and later blocks emphasizing local refinement, which is beneficial for achieving higher final precision in black-box optimization.
 VI-B Deconstructing Component Synergy
-We isolate the contributions of the three core architectural components of L2E: (i) ProxyGrad, which provides gradient injection; (ii) SoftGate, which enables adaptive fusion; and (iii) Mamba, which functions as the structured operator.
-We compare the full model, denoted as Ours, against variants where we individually remove or replace each component.
-VI-B1 Quantitative Impact Analysis
-Table V quantifies the performance degradation that results from the removal of individual components.
-  • Impact of Gradient Guidance (w/o ProxyGrad): The removal of proxy gradients results in the most substantial performance deterioration, as the error increases by $\sim 103\times$.
-This outcome confirms that purely evolutionary search is sample-inefficient in high-dimensional spaces.
-Consequently, proxy gradients provide the descent momentum that is necessary to guide the population.
-  • Impact of Adaptive Fusion (w/o SoftGate): The disablement of the gating mechanism, i.e., the enforcement of a deterministic combination, degrades performance by $\sim 161\times$.
-This result demonstrates the necessity of selective utilization.
-Specifically, the solver must dynamically decide when to trust the gradient for exploitation and when to rely on evolutionary proposals for exploration.
-  • Impact of Architecture (w/o Mamba): The replacement of Mamba with a standard MLP/Attention block increases error by $\sim 6.8\times$.
-Although this degradation is less severe than the removal of gradients, it confirms that Mamba enables the modeling of long-range dependencies and coupled variables.
-This capacity is essential for breaking the precision ceiling on complex landscapes.
-VI-B2 Convergence Dynamics
-Fig.
-11 and Fig.
-12 visualize the per-problem behavior and convergence stability.
-On ill-conditioned functions, such as Ellipsoidal_hc, the Full model (red) achieves the lowest terminal cost and the tightest variance (Fig.
-12, Right).
-In contrast, the w/o SoftGate variant exhibits high variance, which indicates that rigid updates lead to instability.
-The proposed components function in concert.
-Specifically, ProxyGrad provides speed, Mamba ensures structural adaptation to landscape geometry, and SoftGate serves as a robust arbiter.
-The removal of any single component undermines this synergy, thereby leading to slower convergence or stagnation.
-Thus, L2E represents a cohesive system rather than a simple sum of parts.
-Its superior performance stems from the stage-wise curriculum, which is enabled by unshared unrolling, and the adaptive interplay between gradient-based and evolutionary search.
-This interplay is structurally supported by the non-expansive Mamba operator.
+We further isolate the contributions of three core components in L2E: (i) ProxyGrad, which injects proxy numerical guidance; (ii) SoftGate, which enables adaptive fusion; and (iii) Mamba, which parameterizes the structured evolutionary operator.
+We compare the full model (Ours) with variants that remove or replace each component to quantify their individual and combined effects.
 TABLE V: Overall objective on 16 BBOB problems.
@@
 Figure 11: Log-scale objective values (lower is better) on the first eight BBOB problems under four ablation configurations.
-Figure 12: Ablation on convergence and terminal cost across four variants (w/o ProxyGrad, w/o SoftGate, w/o Mamba, and Ours (Full)).
+VI-B1 Quantitative Impact Analysis
+Table V quantifies the performance changes when individual components are removed or replaced.
+  • Impact of Proxy Guidance (w/o ProxyGrad): Removing proxy gradient guidance leads to the largest degradation, increasing the error by $\sim 103\times$.
+This result indicates that, under the considered evaluation budgets, purely evolutionary updates can be substantially less sample-efficient, whereas proxy guidance provides an effective directional signal for accelerating progress.
+  • Impact of Adaptive Fusion (w/o SoftGate): Disabling the gating mechanism (i.e., using a fixed combination) degrades performance by $\sim 161\times$.
+This result underscores the necessity of selective fusion.
+Specifically, the solver benefits from dynamically balancing proxy guidance for refinement and learned evolutionary proposals for exploration.
+In contrast, a rigid mixing rule lacks this essential flexibility.
+    Figure 12: Ablation on convergence and terminal cost across four variants.
 Left: log objective gap vs.
@@
 Right: box-whisker statistics of terminal cost across runs.
+  • Impact of Operator Parameterization (w/o Mamba): Replacing the Mamba-based operator with a standard MLP/attention block increases the error by $\sim 6.8\times$.
+Although this degradation is smaller than that caused by removing guidance or gating, it indicates that the structured operator parameterization enhances the modeling of population-wise interactions and variable couplings.
+This capability is particularly beneficial on complex landscapes.
+VI-B2 Convergence Dynamics
+Figure 11 and Fig.
+12 further visualize convergence behavior and run-to-run variability.
+On ill-conditioned functions such as Ellipsoidal_hc, the full model (red) achieves better terminal performance with tighter variance (Fig.
+12, Right).
+In contrast, the w/o SoftGate variant exhibits noticeably higher variance, consistent with the intuition that rigid update rules can be brittle across diverse instances.
+Overall, the three components act in a complementary manner.
+ProxyGrad provides effective numerical guidance, Mamba offers a structured parameterization for population updates, and SoftGate adaptively mediates between the two update sources.
+Removing any single component weakens this complementarity and often leads to slower progress or earlier stagnation.
+Together with unshared unrolling parameters, these design choices promote stage-dependent behaviors along the trajectory and enable stable, effective optimization across diverse numerical landscapes.
 VII Conclusion
 In this paper, we proposed L2E, a unified framework that approaches evolutionary optimization via neural unrolling.
-L2E enforces strict fixed-point contraction within a bilevel meta-learning scheme.
-This architecture seamlessly integrates the expressiveness of deep learning with the convergence guarantees of classical optimization.
-Empirical results across high-dimensional benchmarks and robot control tasks demonstrate the scalability and generalization capabilities.
-Limitations: However, the reliance on structured neural operators introduces computational overheads.
-Consequently, this computational cost restricts applicability in resource-constrained or low-latency environments.
-Future work will focus on developing lightweight neural operators that are compatible with real-time execution.
-Additionally, we aim to explore sample-efficient first-order algorithms to enhance inference speed.
-Finally, we plan to extend this non-expansive unrolling paradigm to multi-objective and constrained optimization scenarios.
+Instead of learning unconstrained update rules, L2E adopts a stability-inducing unrolling formulation within a bilevel meta-learning scheme, which promotes controlled optimization dynamics in practice.
+Empirical results on high-dimensional numerical benchmarks and robot control tasks demonstrate the scalability of L2E and its robust performance under distribution shifts.
+Limitations: The framework introduces additional training-time overhead due to unrolling and structured operator parameterization.
+This cost may limit applicability in resource-constrained or low-latency settings.
+Future Work: We will develop lightweight operator parameterizations and more efficient training strategies to reduce computation and memory usage.
+We also plan to extend L2E to broader settings, such as constrained and multi-objective optimization.